### v1

# The spectrum of Dirichlet-to-Neumann maps for radial conductivities

Abstract.
The problem of characterizing sequences of real numbers that arise as spectra of Dirichlet-to-Neumann (DtN) maps for elliptic operators has attracted considerable attention over the past fifty years. In this article, we address this question in the simple setting of DtN maps associated with a rotation-invariant elliptic operator $\operatorname{div}(\gamma\nabla\cdot)$ in the ball in Euclidean space. We show that the spectrum of such a DtN operator can be expressed as a universal term, determined solely by the boundary values of the conductivity $\gamma$, plus a sequence of Hausdorff moments of an integrable function, which we call the Born approximation of $\gamma$. We also show that this object is locally determined from the boundary from the corresponding values of the conductivity, a property that implies a local uniqueness result for the Calderón Problem in this setting. We also give a stability result: the functional mapping the Born approximation to its conductivity is Hölder stable in suitable Sobolev spaces. Finally, in order to refine the characterization of the Born approximation, we analyze its regularity properties and their dependence on the conductivity.
1. Introduction
Let $\Omega\subset\mathbb{R}^{d}$, $d\geq 2$, be a smooth bounded domain and $\gamma\in\mathcal{C}(\overline{\Omega})$ a conductivity, i.e. a real-valued continuous function that is bounded below by a strictly positive constant: $\gamma(x)\geq c>0$ for all $x\in\Omega$. The Dirichlet-to-Neumann map on $\partial\Omega$ associated to $\gamma$ is the operator $\Lambda_{\gamma}$ that acts on functions $f\in\mathcal{C}^{\infty}(\partial\Omega)$ as
(1.1) $\Lambda_{\gamma}f:=(\gamma\partial_{\nu}u)|_{\partial\Omega},$
where $\partial_{\nu}$ is the outward normal unit vector field on $\partial\Omega$ and $u$ is the solution to the conductivity equation
(1.2) $\begin{cases}\operatorname{div}(\gamma\nabla u)=0,&\text{ in }\Omega\subset\mathbb{R}^{d},\\ \hfill u=f,&\text{ on }\partial\Omega.\end{cases}$
The operator $\Lambda_{\gamma}$ is an unbounded and positive self-adjoint operator on $L^{2}(\partial\Omega)$ with domain $H^{1}(\partial\Omega)$.
Understanding the structure of the class of all DtN maps $\Lambda_{\gamma}$ as $\gamma$ varies in some class of admissible conductivities is a challenging problem that has multiple implications in various fields: deriving fine properties of their spectrum is a central problem in spectral geometry [CGGS24], and the problem of recovering $\gamma$ from the knowledge of $\Lambda_{\gamma}$ is known as the Calderón inverse problem [FSU25].
In this article, we will present strong structural properties of the set of DtN maps in the presence of rotational symmetry. We will assume that $\Omega=\mathbb{B}^{d}$ is the unit ball in Euclidean space, and we will be interested in conductivities $\gamma$ that are invariant under the action of rotations, that is radial conductivities.
All Dirichlet-to-Neumann operators arising from radial conductivities in $\mathcal{C}(\overline{\mathbb{B}^{d}})$ have the same invariant subspaces, namely, the spaces of spherical harmonics of some fixed degree $k\in\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$. Recall that these are:
[…]
Therefore, denoting for $k\in\mathbb{N}_{0}$
$\Lambda_{\gamma}|_{\mathfrak{H}_{k,d}}=\lambda_{k}[\gamma]\operatorname{\mathrm{Id}}_{\mathfrak{H}_{k,d}},$
one has that $\Lambda_{\gamma}$ is completely determined by its spectrum
$\operatorname{\mathrm{Spec}}(\Lambda_{\gamma})=\{\lambda_{k}[\gamma]\,:\,k\in\mathbb{N}_{0}\}.$
In other words, characterizing radial DtN maps amounts to characterizing their spectra.
Since constants are always solutions to (1.2), one always has $\lambda_{0}[\gamma]=0$. We will show that the rest of the eigenvalues of a radial DtN map have a very rigid structure: they can be expressed in terms of moments of radial functions. For $h\in L^{1}(\mathbb{B}^{d})$ define
(1.3) $\sigma_{k}[h]:=\frac{1}{|\mathbb{S}^{d-1}|}\int_{\mathbb{B}^{d}}h(x)|x|^{2k}\,dx,\qquad k\in\mathbb{N}_{0},$
where $|\mathbb{S}^{d-1}|$ denotes the measure of the unit sphere. The following holds.
Theorem 1.
Let $d\geq 2$, and let $\gamma\in W^{2,p}_{\mathrm{rad}}(\mathbb{B}^{d})$  [^1] Throughout this article, the subscript $\mathrm{rad}$ added to some function space over $\mathbb{B}^{d}$ will mean that we are considering the subspace of its radial elements. with $d/2<p\leq\infty$ be a conductivity. Then there exists a function $\gamma^{\mathrm{B}}\in W^{2,1}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})\subset\mathcal{C}^{1}(\overline{\mathbb{B}^{d}}\setminus\{0\})$ such that
(1.4) $\lambda_{k}[\gamma]=\gamma|_{\mathbb{S}^{d-1}}k-\frac{1}{2}\partial_{\nu}\gamma|_{\mathbb{S}^{d-1}}+\sigma_{k}[\Delta\gamma^{\mathrm{B}}],\qquad\forall k\in\mathbb{N},$
and
(1.5) $\gamma^{\mathrm{B}}|_{\mathbb{S}_{d-1}}=\gamma|_{\mathbb{S}_{d-1}},\qquad\qquad\partial_{\nu}\gamma^{\mathrm{B}}|_{\mathbb{S}_{d-1}}=\partial_{\nu}\gamma|_{\mathbb{S}_{d-1}}.$
In addition, $\gamma^{\mathrm{B}}$ is uniquely determined from $\Lambda_{\gamma}$ since (1.4) and (1.5) imply that $\gamma^{\mathrm{B}}$ is characterized by any of the following equivalent conditions.
  1. i)
    (Solution to a moment problem) $\gamma^{\mathrm{B}}$ is the unique solution in $L^{1}_{\mathrm{rad}}(\mathbb{B}^{d})$ to
    (1.6) $\sigma_{k}[\gamma^{\mathrm{B}}]=\frac{\lambda_{k}[\gamma]}{2(k+1)(k+1+\nu_{d})},\qquad k\in\mathbb{N}_{0},\qquad\nu_{d}:=\frac{d-2}{2}.$
  2. ii)
    (Fourier transform representation) The Fourier transform of $\gamma^{\mathrm{B}}$ is given by
    (1.7) $\widehat{\gamma^{\mathrm{B}}}(\xi):=\int_{\mathbb{B}^{d}}e^{-i\xi\cdot x}\gamma^{\mathrm{B}}(x)\,dx=\pi^{d/2}\sum_{k=1}^{\infty}\frac{(-1)^{k-1}}{k!\Gamma(k+d/2)}\left(\frac{|\xi|}{2}\right)^{2k-2}\lambda_{k}[\gamma],$
    where the series above is absolutely convergent.
The identities (1.4) and (1.7) were already obtained in a formal sense in [BCMM24]. Since $\Delta\gamma^{\mathrm{B}}$ is radial, one can check that the terms $\sigma_{k}[\Delta\gamma^{\mathrm{B}}]$ are actually the Hausdorff moments of a function $h_{\gamma}\in L^{1}([0,1])$:
$\sigma_{k}[\Delta\gamma^{\mathrm{B}}]=\int_{0}^{1}h_{\gamma}(r)r^{k}dr,\qquad k\in\mathbb{N}.$
Sequences of Hausdorff moments have been completely characterized by Hausdorff [Hau23] (see also [Wid41, Chapter 3]); therefore ˜1 provides a partial characterization of the set of radial DtN maps. Condition (1.4) is strictly necessary in order to characterize the spectrum of radial DtN maps: the set of functions $\gamma^{\mathrm{B}}$ obtained as $\gamma$ varies among conductivities in $W^{2,\infty}_{\mathrm{rad}}(\mathbb{B}^{d})$ is strictly contained in $W^{2,1}_{\mathrm{rad}}(\mathbb{B}^{d})$. This follows from the work of Remling [Rem03] on 1-d inverse spectral theory (see Remark˜3.1).
The function $\gamma^{\mathrm{B}}$ appearing in ˜1 is called the Born approximation of $\gamma$. It is a purely spectral object, in the sense that it is completely determined by $\operatorname{\mathrm{Spec}}(\Lambda_{\gamma})$ as (1.7) shows. However, it is local in the sense that in any annular neighborhood of the boundary $\partial\mathbb{B}^{d}$, it is uniquely determined by the values of the conductivity in that same neighborhood.
We define the annulus
(1.8) $U_{s}:=\{x\in\mathbb{R}^{d}\,:\,s<|x|\leq 1\},\qquad 0<s<1.$
Theorem 2 (Uniqueness).
Let $d\geq 2$ and $\gamma_{1},\gamma_{2}\in W^{2,p}_{\mathrm{rad}}(\mathbb{B}^{d})$ with $d/2<p\leq\infty$ be two conductivities. Let $0<s<1$. Then
$\gamma^{\mathrm{B}}_{1}|_{U_{s}}=\gamma^{\mathrm{B}}_{2}|_{U_{s}}\quad\Longleftrightarrow\quad\gamma_{1}|_{U_{s}}=\gamma_{2}|_{U_{s}}.$
These two results refine, in the radial case, structural results for the spectra of DtN maps corresponding to general, smooth, non-radial conductivities obtained, for instance, in [LSA21, Roz78] for the two-dimensional case, or in [PEUVB24] for the ball in Euclidean space, which are based on a detailed analysis of (pseudodifferential or Berezin) symbols of DtN maps. In particular, ˜1 gives a non-trivial description of the spectrum even in the case in which the conductivity $\gamma$ is constant in a neighborhood of the boundary $\partial\mathbb{B}^{d}$. When $\gamma_{1},\gamma_{2}$ are two smooth conductivities that satisfy $\gamma_{1}|_{U_{s}}=\gamma_{2}|_{U_{s}}$ for some $s\in(0,1)$ it is known that $\Lambda_{\gamma_{1}}-\Lambda_{\gamma_{2}}$ is a pseudo-differential operator whose symbol (defined modulo smoothing operators) is zero: this implies that
$\lambda_{k}[\gamma_{1}]-\lambda_{k}[\gamma_{2}]=\mathcal{O}(k^{-\infty}),\qquad k\to\infty.$
A straightforward consequence of Theorems 1 and 2 is the following.
Corollary 3.
Let $d\geq 2$, $d/2<p\leq\infty$, and $s\in(0,1)$. For any two conductivities $\gamma_{1},\gamma_{2}\in W^{2,p}_{\mathrm{rad}}(\mathbb{B}^{d})$ and any $k_{0}\in\mathbb{N}$, the following are equivalent.
  1. i)
    $\gamma_{1}|_{U_{s}}=\gamma_{2}|_{U_{s}}$.
  2. ii)
    There exists $C>0$ such that
    (1.9) $|\lambda_{k}[\gamma_{1}]-\lambda_{k}[\gamma_{2}]|\leq Cs^{2k},\qquad\forall k\geq k_{0}.$
In particular, for every conductivity $\gamma\in W^{2,p}_{\mathrm{rad}}(\mathbb{B}^{d})$,
$\gamma|_{U_{s}}=\gamma|_{\mathbb{S}^{d-1}}\quad\iff\quad\lambda_{k}[\gamma]=\gamma|_{\mathbb{S}^{d-1}}k+\mathcal{O}(s^{2k}),\qquad k\to\infty.$
This result extends in the Euclidean setting [DKN21, Theorem 1.4] to $d\geq 2$ and less regular conductivities. ˜2 has a quantitative counterpart, as we now show.
Define for $N>0$, $K>0$ and $1\leq p\leq\infty$ the set $\mathcal{G}_{K,B}^{p}$ consisting of those conductivities $\gamma\in W^{2,p}(\mathbb{B}^{d})$ that satisfy
(1.10) $K^{-1}\leq\gamma(x)\leq K,\quad x\in\mathbb{B},\qquad\left\lVert\gamma\right\rVert_{W^{2,p}(\mathbb{B}^{d})}\leq N.$
The Born approximation $\gamma^{\mathrm{B}}$ determines a conductivity $\gamma\in\mathcal{G}_{K,B}^{p}\subset{W^{2,1}}(\mathbb{B}^{d})$ in a Hölder continuous way.
Theorem 4 (Stability).
Let $d\geq 2$, $d/2<p<\infty$, $K>1$. There exist $N_{0}>0$ such that, for $N>N_{0}$, the following assertions hold:
  1. i)
    For all $0<s<1$ there exist $0<\delta<1$ and $C_{N,K,d,p,s}>0$ such that, for every $\gamma_{1},\gamma_{2}\in\mathcal{G}_{K,N}^{p}$ satisfying $\left\lVert\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2}\right\rVert_{W^{2,1}(\mathbb{B}^{d})}<\delta$, one has
    $\left\lVert\gamma_{1}-\gamma_{2}\right\rVert_{W^{2,1}(U_{s})}\leq C_{N,K,d,p,s}\left\lVert\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2}\right\rVert_{W^{2,1}(U_{s})}^{\frac{p-1}{2p-1}}.$
  2. ii)
    For all $1<q<p$, there exist $0<\delta<1$, $0<\alpha<1$ and $C_{N,K,d,p,q}>0$ such that for every $\gamma_{1},\gamma_{2}\in\mathcal{G}_{K,N}^{p}$ satisfying $\left\lVert\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2}\right\rVert_{W^{2,1}(\mathbb{B}^{d})}<\delta$ one has
    $\left\lVert\gamma_{1}-\gamma_{2}\right\rVert_{W^{2,q}(\mathbb{B}^{d})}\leq C_{N,K,d,p,q}\left\lVert\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2}\right\rVert_{W^{2,1}(\mathbb{B}^{d})}^{\alpha}.$
The stability estimate i) shows that the Hölder modulus of continuity of the map $\gamma^{\mathrm{B}}\mapsto\gamma$ only depends on $p$ in the annuli $U_{s}$. This local estimate can be extended to the global estimate ii) on $\mathbb{B}^{d}$ if one allows the Hölder modulus of continuity to depend on $N,K,p,d$. By interpolation, one can also get the norm $W^{2,q}(\mathbb{B}^{d})$ in the LHS of the estimate i) paying the price of a worse Hölder modulus of continuity in the RHS.
˜4 shows that $\gamma$ can be determined by $\gamma^{\mathrm{B}}$ in a Hölder stable way, which is not the case for the problem of recovering a radial $\gamma$ from $\operatorname{\mathrm{Spec}}(\Lambda_{\gamma})$: this problem is ill-posed (no continuity unless the conductivities lie in a compact set [Ale88, AC08, FKR14] and the modulus of continuity is generally, at best, logarithmic [Man01]). Therefore, this shows that the ill-posedness of reconstructing a radial $\gamma$ from $\operatorname{\mathrm{Spec}}(\Lambda_{\gamma})$ is due to the linear step of solving the moment problem (1.4) or (1.6) to reconstruct $\gamma^{\mathrm{B}}$ (the moment problem is, indeed, a notoriously ill-posed problem). Thus, the introduction of the Born approximation $\gamma^{\mathrm{B}}$, which depends linearly on $\operatorname{\mathrm{Spec}}(\Lambda_{\gamma})$, allows transforming an ill-posed problem into a Hölder stable one.
One can compute the Born approximation of a conductivity explicitly in some special cases, see Section˜3.5; some of these explicit examples are shown in Fig.˜1. Also, it is possible to infer more precise information on the regularity of the Born approximation. In general, it is always a continuous function on $\mathbb{B}^{d}\setminus\{0\}$ that can be eventually singular at the origin. The nature of this singularity is related to the existence of eigenvalues and/or resonances for a certain 1-d Schrödinger operator defined in terms of $\gamma$, as will be shown in Section˜3.
Theorem 5 (Regularity).
Let $d\geq 2$ and let $\gamma\in W^{2,p}_{\mathrm{rad}}(\mathbb{B}^{d})$ with $d/2<p\leq\infty$ be a conductivity. If $\nu_{d}:=(d-2)/2$, the following assertions hold.
  1. i)
    If $d>2$, there exists a finite set of real numbers $0<\kappa_{1}<\dots<\kappa_{J}<\nu_{d}$, $J\in\mathbb{N}_{0}$, (possibly empty) and constants $c_{0}\in\mathbb{R},\,c_{1},\cdots,c_{J}\geq 0$ such that
    $\gamma^{\mathrm{B}}(x)=-c_{0}\log|x|+\sum_{j=1}^{J}\frac{c_{j}}{|x|^{2\kappa_{j}}}+\gamma^{\mathrm{B}}_{\mathrm{reg}}(x),$
    where $\gamma^{\mathrm{B}}_{\mathrm{reg}}\in\mathcal{C}(\overline{\mathbb{B}^{d}})\cap W^{1,d}(\mathbb{B}^{d})\cap W^{2,1}_{\mathrm{rad}}(\mathbb{B}^{d})$.
  2. ii)
    If $d=2$ then $\gamma^{\mathrm{B}}\in\mathcal{C}(\overline{\mathbb{B}^{2}})\cap W^{2,1}_{\mathrm{rad}}(\mathbb{B}^{2})$.
  3. iii)
    $\gamma^{\mathrm{B}}|_{\mathbb{B}^{d}\setminus\{0\}}\in W^{2,p}_{\mathrm{loc}}(\mathbb{B}^{d}\setminus\{0\})$ and for every $m\geq 2$ and $0<s<1$ one has $\gamma^{\mathrm{B}}\in\mathcal{C}^{m}(U_{s})$ if and only if $\gamma\in\mathcal{C}^{m}(U_{s})$.
The notion of Born approximation used in this work was introduced in [BCMM22] and [BCMM24] as a formal object satisfying (1.6) or (1.7), in the context of the Calderón Problem. This problem goes back to Calderón [Cal80], [^2] This reference has been reprinted in [Cal06]. who considered it during the fifties, and has been the subject of intense research in the last forty years, partly because of its applications to medical imaging techniques, such as Electrical Impedance Tomography. We refer the reader to the recent book [FSU25] in order to obtain a wider perspective on this problem as well as an up-to-date guide to the vast literature on the subject.
From this point of view, the existence of the Born approximation proved in ˜1 provides a decomposition of the process of reconstructing $\gamma$ from $\Lambda_{\gamma}$ in two steps.
  • The first step (obtaining $\gamma^{\mathrm{B}}$ from $\Lambda_{\gamma}$) is linear and amounts to solving the moment problem (1.6), which is a notoriously ill-posed inverse problem.
  • The second step (obtaining $\gamma$ from $\gamma^{\mathrm{B}}$) is non-linear but enjoys good stability properties, as ˜4 shows.
The second property shows that the Born approximation induces a stable factorization of the Calderón problem in the radial case. Also, the local nature of $\gamma^{\mathrm{B}}$ provides a direct way to obtain local uniqueness for the radial Calderón problem (˜3) without using the standard approach based on Complex Geometrical Optics solutions of the equation (1.2). This approach based on the Born approximation has also been successfully applied to the Schrödinger operator version of the Calderón problem (the so-called Gelf’and-Calderón problem) at zero energy [DMMN24], and at fixed energy [MMSM25], and has been formulated in more general settings in [MM24]. We remark that, using the reconstruction algorithms in [DMMN24], one could transform ˜3 into an explicit reconstruction algorithm to determine $\gamma$ from $\Lambda_{\gamma}$.
The proofs of the previous theorems use the notion of $A$-amplitude introduced by Simon in [Sim99] in the context of inverse spectral theory for half-line Schrödinger operators. This notion of $A$-amplitude, further developed in [GS00, RS00], is also a fundamental tool in [DMMN24, MMSM25]. In particular, to prove ˜4 we use stability results for the problem of reconstructing a Schrödinger potential from its $A$-amplitude developed in [DMMN24, Section 5]. The connection between the $A$-amplitude and $\gamma^{\mathrm{B}}$ is not superficial: it has been shown in [MM24] that the $A$-amplitude is a Born approximation for the inverse problem of reconstructing a Schrödinger potential from its Weyl-Titchmarsh function. The $A$-amplitude also plays an important role in the closely related problem of studying spectra of DtN maps in warped-product manifolds with boundary, see [DKN20, DKN21, DKN24].
Structure of the article.
˜1, ˜2, ˜3 and ˜5 are proved in Section˜2, whereas ˜4 is proved in Section˜4. Section˜3 contains a detailed analysis of the structure of singularities of the Born approximation, both in the Schrödinger operator and conductivity cases, that are used in the proofs of ˜1 and ˜5; these results are of independent interest.
Acknowledgments
F.M. and C.M. have been supported by grant PID2021-124195NB-C31 from Agencia Estatal de Investigación (Spain). F. N. is supported by the GDR Dynqua.
2. Existence of the Born approximation
2.1. Linearized DtN operators in the radial case
Consider the non-linear map
(2.1) $\begin{array}[]{cccl}\Psi:&X&\longrightarrow&\mathcal{L}(H^{1/2}(\partial\mathbb{B}^{d}),H^{-1/2}(\partial\mathbb{B}^{d}))\\ &\gamma&\longmapsto&\Lambda_{\gamma}\end{array}$
where $X$ denotes the set of conductivities that are in $L^{\infty}(\mathbb{B}^{d})$. In this section, we obtain an explicit formula for the Fréchet differential in the radial case.
Proposition 2.1.
The Fréchet derivative $d\Psi_{\mathds{1}}$ of $\Psi$ in the space $L^{\infty}(\mathbb{B}^{d},\mathbb{R})$ at the conductivity $\gamma=\mathds{1}$ exists and it is a bounded operator
$d\Psi_{\mathds{1}}:L^{\infty}(\mathbb{B}^{d},\mathbb{R})\to\mathcal{L}(H^{1}(\mathbb{S}^{d-1}),L^{2}(\mathbb{S}^{d-1})).$
If, in addition, $h\in L^{\infty}_{\mathrm{rad}}(\mathbb{B}^{d})$, then $d\Psi_{\mathds{1}}(h)$ is the unbounded operator on $L^{2}(\mathbb{S}^{d-1})$, with domain $H^{1}(\mathbb{S}^{d-1})$, characterized by
(2.2) $d\Psi_{\mathds{1}}(h)|_{\mathfrak{H}_{k}}=k\left(2k+d-2\right)\sigma_{k-1}[h]\operatorname{\mathrm{Id}}_{\mathfrak{H}_{k}},\qquad k\in\mathbb{N},$
and $d\Psi_{\mathds{1}}(h)(1)=0$. In particular $\Lambda_{\mathds{1}}=d\Psi_{\mathds{1}}(\mathds{1})$, and $d\Psi_{\mathds{1}}$ admits a natural extension to $L^{1}_{\mathrm{rad}}(\mathbb{B}^{d})$ via identity (2.2).
Proof of Proposition˜2.1.
The Fréchet differential of $\Psi:\gamma\longmapsto\Lambda_{\gamma}$ in $L^{\infty}_{\mathrm{rad}}(\mathbb{B}^{d})$ at the constant conductivity $\gamma=1$ is given by the expression
$(f_{1},d\Psi_{\mathds{1}}(h)f_{2})_{L^{2}(\mathbb{S}^{d-1})}=\int_{\mathbb{B}^{d}}h(x)\overline{\nabla u_{1}(x)}\cdot\nabla u_{2}(x)\,dx,$
where $\Delta u_{j}=0$ in $\mathbb{B}^{d}$ and $u_{j}|_{\mathbb{S}^{d-1}}=f_{j}$, for $j=1,2$. Therefore, $d\Psi_{\mathds{1}}(h)$ is diagonalizable on the basis of spherical harmonics and its eigenvalues are:
(2.3) $(Y_{k},d\Psi_{\mathds{1}}(h)Y_{k})_{L^{2}(\mathbb{S}^{d-1})}=\int_{\mathbb{B}^{d}}h(x)\overline{\nabla Y_{k}^{s}(x)}\cdot\nabla Y_{k}^{s}(x)\,dx,$
where $Y_{k}^{s}(x):=|x|^{k}Y_{k}(x/|x|)$ stands for the solid harmonic corresponding to a normalized spherical harmonic $Y_{k}\in\mathfrak{H}_{k}$. When $k=0$, the solid harmonics are constant functions and (2.3) implies $d\Psi_{\mathds{1}}(h)(1)=0$.
(a) $d=2,\;\mu=1,\;\nu=3$.
(b) $d=3,\;\mu=1,\;\nu=3$.
(c) $d=2,\;\mu=3,\;\nu=1$.
(d) $d=3,\;\mu=3,\;\nu=1$.
(e) $d=3,\;\mu=1,\;\nu=0$.
Figure 1. Plots of the radial profiles of $\gamma_{d,\mu,\nu}$ (blue) and $\gamma^{\mathrm{B}}_{d,\mu,\nu}$ (orange) defined in Section˜3.5.
Direct computation gives that
$\nabla Y_{k}^{s}(x)=|x|^{k-1}\left(kY_{k}(\hat{x})\hat{x}+\nabla_{\mathbb{S}^{d-1}}Y_{k}(\hat{x})\right),\qquad\hat{x}=x/|x|.$
where $\nabla_{\mathbb{S}^{d-1}}$ denotes the gradient in $\mathbb{S}^{d-1}$. Writing $h=h_{0}(|\cdot|)$ and using that $\Delta_{\mathbb{S}^{d-1}}Y_{k}=k(k+d-2)Y_{k}$. $\left\lVert Y_{k}\right\rVert_{L^{2}(\mathbb{S}^{d-1})}=1$, one gets that (2.3) equals
$\left(k^{2}+\int_{\mathbb{S}^{d-1}}|\nabla_{\mathbb{S}^{d-1}}Y_{k}(x)|^{2}\,dS(x)\right)\int_{0}^{1}h_{0}(r)r^{2k-2+d-1}\,dr=k(2k-d-2)\sigma_{k-1}[h],$
as claimed. In the particular case $h=\mathds{1}$ this gives
$(Y_{k},d\Psi_{\mathds{1}}(\mathds{1})Y_{k})_{L^{2}(\mathbb{S}^{d-1})}=k$
which implies that $d\Psi_{\mathds{1}}(\mathds{1})=\Lambda_{\mathds{1}}$. ∎
2.2. The Born approximation for Schrödinger operator DtN maps
It is well-known that, provided $\gamma$ has enough regularity, the study of $\Lambda_{\gamma}$ can be reduced to that of a DtN map associated to a certain Schrödinger operator. If $\gamma\in W^{2,p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R}_{+})$ with $p>d/2$ is a conductivity, and $u$ solves (1.2) then $v=\sqrt{\gamma}u$ satisfies
(2.4) $\begin{cases}-\Delta v+Vv&=0,\qquad\text{on }\mathbb{B}^{d},\\ \hfill v|_{\mathbb{S}^{d-1}}&=\sqrt{\gamma}f,\end{cases}$
where
(2.5) $V:=\frac{\Delta\sqrt{\gamma}}{\sqrt{\gamma}}\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R}).$
One can check that the following identity holds
(2.6) $\Lambda^{\mathrm{p}}_{V}f:=\partial_{\nu}v|_{\mathbb{S}^{d-1}}=\gamma^{-1/2}\Lambda_{\gamma}(\gamma^{-1/2}f)+\frac{1}{2}\gamma^{-1}(\partial_{\nu}\gamma)f,$
when $v$ solves (2.4) with $V$ given by (2.5). From this identity, it is clear that the Calderón problem for conductivities is equivalent to the one with Schrödinger operators provided $\gamma$ and $\partial_{\nu}\gamma$ are known at the boundary $\mathbb{S}^{d-1}$.
Again, since (2.4) is invariant by rotations,
(2.7) $\Lambda^{\mathrm{p}}_{V}|_{\mathfrak{H}_{k}}=\lambda^{\mathrm{p}}_{k}[V]\operatorname{\mathrm{Id}}_{\mathfrak{H}_{k}},\qquad\forall k\in\mathbb{N}_{0}.$
Moreover, when $V=0$ a direct computation gives that
$\Lambda^{\mathrm{p}}_{0}(Y_{k})=kY_{k}\qquad\forall Y_{k}\in\mathfrak{H}_{k},$
and hence, $\lambda^{\mathrm{p}}_{k}[0]=\lambda_{k}[\mathds{1}]=k$ for every $k\in\mathbb{N}_{0}$, since one has $\Lambda^{\mathrm{p}}_{0}=\Lambda_{\mathds{1}}$ from (2.6).
The following result improves the conclusion of [DMMN24, Theorem 1] when the potential arises from a conductivity via identity (2.5). For shortness, define
$\left\lVert f\right\rVert_{d,\alpha}:=\int_{\mathbb{B}^{d}}|f(x)||x|^{2-d}(1-\log|x|)^{\alpha}\,dx<\infty,\qquad\alpha\geq 0,\,d\geq 2.$
Proposition 2.2.
Let $d\geq 2$ and let $\gamma\in W^{2,p}_{\mathrm{rad}}(\mathbb{B}^{d})$ with $d/2<p\leq\infty$ be a conductivity. Let $V$ be given by (2.5) and $\Lambda^{\mathrm{p}}_{V}$ by (2.6). Then $V\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$ and there exists a unique $V^{\mathrm{B}}\in L^{1}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$ such that
(2.8) $\sigma_{k}[V^{\mathrm{B}}]=\lambda_{k}[V]-k,\qquad\text{for all }\,k\in\mathbb{N}_{0}.$
In addition:
  1. i)
    If $d>2$, there exists a finite (possibly empty) set of real numbers $0<\kappa_{1}<\dots<\kappa_{J}<\nu_{d}$ with $J\in\mathbb{N}_{0}$ and some constants $a_{0}\in\mathbb{R},\,a_{1},\dots,a_{J}\leq 0$ such that
    (2.9) $V^{\mathrm{B}}(x)=\frac{a_{0}}{|x|^{2}}+\sum_{j=1}^{J}\frac{a_{j}}{|x|^{2\kappa_{j}+2}}+V^{\mathrm{B}}_{\mathrm{reg}}(x),$
    where $V^{\mathrm{B}}_{\mathrm{reg}}$ satisfies $\left\lVert V^{\mathrm{B}}_{\mathrm{reg}}\right\rVert_{d,\alpha}<\infty$ for all $\alpha>1$.
  2. ii)
    If $d=2$ then $\left\lVert V^{\mathrm{B}}\right\rVert_{2,\alpha}<\infty$ for all $\alpha>1$.
  3. iii)
    $V^{\mathrm{B}}|_{\mathbb{B}^{d}\setminus\{0\}}\in L^{p}_{\mathrm{loc}}(\mathbb{B}^{d}\setminus\{0\})$. In fact $V^{\mathrm{B}}-V$ is continuous on $\overline{\mathbb{B}^{d}}\setminus\{0\}$ and vanishes in $\partial\mathbb{B}^{d}$. Also, $V\in\mathcal{C}^{m}(U_{s})$ iff $V^{\mathrm{B}}\in\mathcal{C}^{m}(U_{s})$, for all $m\in\mathbb{N}_{0}$ and for all $0<s<1$.
Proof of Proposition˜2.2.
If $V$ is given by (2.5), then $V\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$, since $\gamma(x)>c>0$ in $\mathbb{B}^{d}$. From Lemma˜3.9 and Proposition˜3.10 i) it follows there exists a solution $V^{\mathrm{B}}$ of the moment problem (2.8). This solution is unique as follows from [BCMM22, Equation 1.20], which is reproduced here in (2.14) below. Then i) and ii) are immediate from Lemma˜3.9 and Proposition˜3.10. Finally, assertion iii) is proved in [DMMN24, Theorem 5]. ∎
2.3. The Born approximation for $\Lambda_{\gamma}$
The proof of ˜1 follows from Proposition˜2.2, Proposition˜2.1 and the following technical properties of radial functions.
Lemma 2.3.
Let $d\geq 2$. The following assertions hold:
  1. i)
    If $u\in W^{2,1}_{\mathrm{rad}}(\mathbb{B}^{d})$ then $u\in\mathcal{C}^{1}(\overline{\mathbb{B}^{d}}\setminus\{0\})$. Moreover, if $u=u_{0}(|\cdot|)$, the following quantities are well-defined
    (2.10) $a(u):=u_{0}(1),\qquad b(u):=\frac{1}{2}\partial_{r}u_{0}(1),$
    and satisfy $|a(u)|\leq C\left\lVert u\right\rVert_{W^{2,1}(\mathbb{B}^{d})}$, $|b(u)|\leq C\left\lVert u\right\rVert_{W^{2,1}(\mathbb{B}^{d})}$ for some $C>0$.
  2. ii)
    For all $u\in W^{2,1}_{\mathrm{rad}}(\mathbb{B}^{d})$
    (2.11) $\sigma_{0}[\Delta u]$ $=2b(u),$
    (2.12) $\sigma_{k}[\Delta u]$ $=2b(u)-2ka(u)+2k(2k+d-2)\sigma_{k-1}[u],\qquad\forall k\in\mathbb{N}.$
  3. iii)
    For all $u\in L^{1}_{\mathrm{rad}}(\mathbb{B}^{d})$
    (2.13) $\widehat{u}(\xi)=\pi^{d/2}\sum_{k=1}^{\infty}\frac{(-1)^{k-1}}{k!\Gamma(k+d/2)}\left(\frac{|\xi|}{2}\right)^{2k-2}k(2k+d-2)\sigma_{k-1}[u].$
Proof.
Statement i) is immediate using that $\partial_{r}^{2}u_{0}(r)$ belongs to $L^{1}(\varepsilon,1)$ for all $\varepsilon>0$. The boundedness of the trace functionals $a$ and $b$ follows immediately by the standard trace theorem with exponent $p=1$. We now prove the remaining items. By integration by parts:
$\sigma_{k}[\Delta u]=\int_{0}^{1}$ $\partial_{r}(r^{d-1}\partial_{r}u_{0}(r))r^{2k}\,dr=\partial_{r}u_{0}(1)-2k\int_{0}^{1}\partial_{r}u_{0}(r)r^{2k+d-2}\,dr$
$=\partial_{r}u_{0}(1)-2ku_{0}(1)+2k(2k+d-2)\int_{0}^{1}u_{0}(r)r^{2k+d-3}\,dr$
$=\partial_{r}u_{0}(1)-2ku_{0}(1)+2k(2k+d-2)\sigma_{k-1}[u].$
Thus, for any $u\in W^{2,1}_{\mathrm{rad}}(\mathbb{B}^{d})$ we have
$\sigma_{k}[\Delta u]=\partial_{r}u_{0}(1)-2ku_{0}(1)+2k(2k+d-2)\sigma_{k-1}[u],$
which proves (2.12). The previous derivation also holds with $k=0$, in which case one gets $\sigma_{0}[\Delta u]=\partial_{r}u_{0}(1)$. This concludes the proof of ii). Formula (2.13) follows from [BCMM22, Equation 1.20], which states that
(2.14) $\widehat{u}(\xi)=2\pi^{d/2}\sum_{k=0}^{\infty}\frac{(-1)^{k}}{k!\Gamma(k+d/2)}\left(\frac{|\xi|}{2}\right)^{2k}\sigma_{k}[u],$
and straightforward algebraic properties of the gamma function. ∎
Define the following operator acting on $L^{1}_{\mathrm{rad}}(\mathbb{B}^{d})$.
(2.15) $T_{d}f(x):=\int_{|x|}^{1}\frac{1}{t^{d-1}}\int_{0}^{t}f_{0}(s)s^{d-1}\,dsdt,\qquad f\in L^{1}_{\mathrm{rad}}(\mathbb{B}^{d}),\;f=f_{0}(|\cdot|).$
One readily checks that $T_{d}f\in\mathcal{C}^{1}(\overline{\mathbb{B}^{d}}\setminus\{0\})$ is radial, $T_{d}f|_{\mathbb{S}^{d-1}}=0$, and that $\Delta T_{d}f=f$ a.e. in $\mathbb{B}^{d}$.
Lemma 2.4.
Let $f\in L^{1}_{\mathrm{rad}}(\mathbb{B}^{d})$ and $u:=T_{d}(f)$.
  1. i)
    If $f\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d})$ for some $1<p<\infty$, then $u\in W^{2,p}_{\mathrm{rad}}(\mathbb{B}^{d})$.
  2. ii)
    If $f$ satisfies for some $\alpha>1$ that
    $\left\lVert f\right\rVert_{d,\alpha}=\int_{\mathbb{B}^{d}}|f(x)||x|^{2-d}(1-\log|x|)^{\alpha}\,dx<\infty,$
    then $u\in\mathcal{C}(\overline{\mathbb{B}^{d}})$, $\nabla u\in L^{d}(\mathbb{B}^{d})$, and $D^{2}u\in L^{1}(\mathbb{B}^{d})$.
Proof.
i) Since $\Delta u=f$, $f\in L^{p}(\mathbb{B}^{d})$ and $u$ vanishes on $\mathbb{S}^{d-1}$, $u\in W^{2,p}(\mathbb{B}^{d})$ by the standard Calderon-Zygmund estimates for the Laplacian.
ii) Let $F$ denote the radial profile of $u$. For all $\alpha>1$, we have the estimate
$|F^{\prime}(r)|\leq\frac{1}{r}\int_{0}^{r}|f_{0}(s)|s\,ds\leq\frac{1}{r(1-\log r)^{\alpha}}\int_{0}^{r}|f_{0}(s)|s(1-\log s)^{\alpha}\,ds\leq\frac{C_{\alpha}}{r(1-\log r)^{\alpha}}.$
This implies that $F(0)$ is well defined and finite, since, under our hypothesis on $\alpha$,
$|F(0)|\leq\int_{0}^{1}|F^{\prime}(r)|\,dr\leq\int_{0}^{1}\frac{1}{r(1-\log r)^{\alpha}}\,dr=\int_{0}^{\infty}\frac{1}{(1+t)^{\alpha}}dt<\infty.$
The integrability of $F^{\prime}$ in $[0,1]$ implies that $F$ is continuous on $[0,1]$, and hence, $u\in\mathcal{C}(\overline{\mathbb{B}^{d}})$. Since $\alpha>1>1/d$, one finds that the gradient satisfies
$\left\lVert\nabla u\right\rVert_{L^{d}(\mathbb{B}^{d})}^{d}\leq|\mathbb{S}^{d-1}|\int_{0}^{1}|F^{\prime}(r)|^{d}r^{d-1}\,dr\leq|\mathbb{S}^{d-1}|\int_{0}^{1}\frac{1}{r(1-\log r)^{d\alpha}}\,dr<\infty.$
One can similarly bound the Hessian, using $F^{\prime\prime}(r)=-(d-1)r^{-1}F^{\prime}(r)-f_{0}(r)$,
$|\partial_{x_{i}x_{j}}^{2}u(x)|\leq|F^{\prime\prime}(|x|)|+2|F^{\prime}(|x|)||x|^{-1}\leq(d+1)|F^{\prime}(|x|)||x|^{-1}+|f(x)|.$
and,
$\left\lVert D^{2}u\right\rVert_{L^{p}(\mathbb{B}^{d})}\leq(d+1)\left(\int_{\mathbb{B}^{d}}|F^{\prime}(|x|)|^{p}|x|^{d-1-p}\,dx\right)^{1/p}+\left\lVert f\right\rVert_{L^{p}(\mathbb{B}^{d})}\\ \leq C_{d,\alpha}\left(\int_{0}^{1}\frac{r^{d-1-2p}}{(1-\log r)^{p\alpha}}\,dr\right)^{1/p}+\left\lVert f\right\rVert_{L^{p}(\mathbb{B}^{d})}.$
The first integral converges whenever $1\leq p\leq d/2$, $\alpha>1$, and, by hypothesis, $\left\lVert f\right\rVert_{L^{1}(\mathbb{B}^{d})}<\infty$; therefore $\left\lVert D^{2}u\right\rVert_{L^{1}(\mathbb{B}^{d})}<\infty$ as we wanted to prove. ∎
Proof of ˜1.
Let $V$ be given by (2.4) and restrict the operator identity (2.6) to functions in $\mathfrak{H}_{k}$ to obtain
(2.16) $\lambda^{\mathrm{p}}_{k}[V]-k=\frac{1}{a(\gamma)}\left(\lambda_{k}[\gamma]+b(\gamma)-a(\gamma)k\right),\qquad\forall\,k\in\mathbb{N}_{0},$
where the notation in (2.10) has been used. On the other hand, since $\gamma\in W^{2,p}_{\mathrm{rad}}(\mathbb{B}^{d})$, we have $V\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$. Therefore, Proposition˜2.2 applies and there exists $V^{\mathrm{B}}\in L^{1}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$ such that
$\lambda^{\mathrm{p}}_{k}[V]-k=\sigma_{k}[V^{\mathrm{B}}],\qquad\forall\,k\in\mathbb{N}_{0}.$
Inserting this in (2.16) yields
(2.17) $\lambda_{k}[\gamma]=a(\gamma)k-b(\gamma)+a(\gamma)\sigma_{k}[V^{\mathrm{B}}],\qquad\forall\,k\in\mathbb{N}_{0}.$
Define
(2.18) $\gamma^{\mathrm{B}}(x):=a(\gamma)-2a(\gamma)T_{d}(V^{\mathrm{B}}).$
Then $\gamma^{\mathrm{B}}\in\mathcal{C}^{1}(\overline{\mathbb{B}^{d}}\setminus\{0\})$. By Proposition˜2.2, $V^{\mathrm{B}}=V^{\mathrm{B}}_{\mathrm{sing}}+V^{\mathrm{B}}_{\mathrm{reg}}$ where $\left\lVert V^{\mathrm{B}}_{\mathrm{reg}}\right\rVert_{d,\alpha}<\infty$ for all $\alpha>1$, and either $V^{\mathrm{B}}_{\mathrm{sing}}=0$, if $d=2$, or $V^{\mathrm{B}}_{\mathrm{sing}}\in L^{q}_{\mathrm{rad}}(\mathbb{B}^{d})$ for every $1\leq q<d/s_{V}$, with $s_{V}=2\max_{j=0,...,J}\{1,\kappa_{j}+1\}<d$ when $d\geq 3$. Therefore, by Lemma˜2.4, one gets that $\gamma^{\mathrm{B}}\in W^{2,1}_{\mathrm{rad}}(\mathbb{B}^{d})$. Also, $V^{\mathrm{B}}$ solves the Poisson problem
(2.19) $\Delta\gamma^{\mathrm{B}}=2a(\gamma)V^{\mathrm{B}},\qquad a(\gamma^{\mathrm{B}})=a(\gamma).$
This proves (1.4) by direct substitution in (2.17).
To prove (1.5), it remains to show that $b(\gamma)=b(\gamma^{\mathrm{B}})$. Evaluating (2.17) in $k=0$ and using that $\lambda_{0}[\gamma]=0$, we get that $b(\gamma)=a(\gamma)\sigma_{0}[V^{\mathrm{B}}]$. Using that $\Delta\gamma^{\mathrm{B}}=2a(\gamma)V^{\mathrm{B}}$ and (2.11), we obtain
$b(\gamma)=a(\gamma)\sigma_{0}[V^{\mathrm{B}}]=\frac{1}{2}\sigma_{0}[\Delta\gamma^{\mathrm{B}}]=b(\gamma^{\mathrm{B}}).$
Hence the identities (1.4) and (1.5) hold. This proves that the Born approximation exists.
Applying (2.12) to $\gamma^{\mathrm{B}}$ and inserting the identity (1.5) gives
$\frac{1}{2}\sigma_{k}[\Delta\gamma^{\mathrm{B}}]=b(\gamma)-ka(\gamma)+k(2k+d-2)\sigma_{k-1}[\gamma^{\mathrm{B}}],\quad\forall k\in\mathbb{N}.$
Combining this identity with (1.4) we deduce that
$\lambda_{k}[\gamma]=k\left(2k+d-2\right)\sigma_{k-1}[\gamma^{\mathrm{B}}],\qquad\forall k\in\mathbb{N},$
or, equivalently, that $\gamma^{\mathrm{B}}$ is a solution of the moment problem (1.6). This problem has uniqueness of solutions in $L^{1}_{\mathrm{rad}}(\mathbb{B}^{d})$, as the reconstruction formula (2.13) implies (in fact, from [DMMN24, Lemma 6.2] it follows that uniqueness holds even for compactly supported, radial distributions). Identity (1.7) follows directly from (2.13). ∎
Proof of ˜2.
Assume first that $\gamma^{\mathrm{B}}_{1}=\gamma^{\mathrm{B}}_{2}$ on $U_{s}$. Then, in particular (1.5) implies
(2.20) $\gamma_{1}|_{\mathbb{S}_{d-1}}=\gamma_{2}|_{\mathbb{S}_{d-1}},\qquad\qquad\partial_{\nu}\gamma_{1}|_{\mathbb{S}_{d-1}}=\partial_{\nu}\gamma_{2}|_{\mathbb{S}_{d-1}}.$
Also, we get $V^{\mathrm{B}}_{1}=V^{\mathrm{B}}_{2}$ on $U_{s}$ by (2.19), which implies $V_{1}=V_{2}$ on $U_{s}$ by [DMMN24, Theorem 2]. Then, if $u:=\sqrt{\gamma_{1}}-\sqrt{\gamma_{2}}$, in $U_{s}$ we have by (2.4) and (2.20) that
[…]
where $V:=V_{1}=V_{2}$ on $U_{s}$. Since $V\in L^{p}(U_{s})$ with $p>\frac{d}{2}$, by unique continuation, we deduce that $u=0$ on $U_{s}$ and therefore $\gamma_{1}=\gamma_{2}$ on $U_{s}$.
If we now assume $\gamma_{1}=\gamma_{2}$ on $U_{s}$, then (2.4) implies $V_{1}=V_{2}$ on $U_{s}$, and hence one gets $V^{\mathrm{B}}_{1}=V^{\mathrm{B}}_{2}$ on $U_{s}$ by [DMMN24, Theorem 2]. Finally, (2.19) and (2.20) (which again holds by (1.5)) imply that $\gamma^{\mathrm{B}}_{1}=\gamma^{\mathrm{B}}_{2}$ on $U_{s}$ by unique continuation. ∎
Proof of ˜3.
This result is a straightforward consequence of ˜1, ˜2 and the following fact.
Lemma 2.5.
For $F\in L^{1}_{\mathrm{rad}}(\mathbb{B}^{d})$, $s\in(0,1)$, and $k_{0}\in\mathbb{N}$, the following holds:
$F|_{U_{s}}=0\;\iff\;\exists C>0,\quad|\sigma_{k}[F]|\leq Cs^{2k},\quad k\geq k_{0}.$
Proof.
One implication is straightforward; in order to check the other, assume that $|\sigma_{k}[F]|\leq Cs^{2k}$ for $k\geq k_{0}$. If $F=f(|\cdot|)$ one has $\sigma_{k}[F]=\mathcal{L}(F)(k+d/2)$ where
$\mathcal{L}(F)(z):=\int_{0}^{\infty}f(e^{-t})e^{-z2t}dt=\frac{1}{2}\int_{0}^{\infty}f(e^{-t/2})e^{-zt}dt.$
By hypothesis, $e^{(2k+d)|\log s|}|\mathcal{L}(F)(k+d/2)|\leq Ce^{-d|\log s|}$ for every $k\geq k_{0}$. Note that there exists $C_{s}>0$ such that
(2.21) $\left|\int_{0}^{2|\log s|}f(e^{-t/2})e^{-zt}dt\right|\leq|\mathcal{L}(F)(z)|+C_{s}\left\lVert F\right\rVert_{L^{1}(\mathbb{B}^{d})}e^{-z2|\log s|},\qquad z>d/2.$
Define
$G(z):=e^{z2|\log s|}\int_{0}^{2|\log s|}f(e^{-t/2})e^{-zt}dt,$
we have, in view of (2.21), that $G$ is an entire function that satisfies
$\sup_{k\in\mathbb{N}}|G(k+d/2)|\leq Ce^{-d|\log s|}+C_{s}\left\lVert F\right\rVert_{L^{1}(\mathbb{B}^{d})}$
and
$|G(z)|\leq\left(\int_{0}^{2|\log s|}|f(e^{-t})|dt\right)e^{2|\log s|(\operatorname{\mathrm{Re}}z)_{+}}.$
One can apply a theorem of Duffin and Schaeffer [Boa54, Theorem 10.5.1] to conclude that $M>0$ exists such that $|G(z)|\leq M$ when $z>0$. This implies that
$\int_{0}^{2|\log s|}f(e^{-t/2})e^{-zt}dt=\mathcal{O}(e^{-z2|\log s|}),\qquad z\to\infty.$
Using [Sim99, Lemma A.2.1] we infer that $f(t)=0$ for a.e. $t\in(0,s)$, and the conclusion follows. ∎
To see that two conductivities satisfying (1.9) must necessarily coincide on $U_{s}$, start noticing that (1.4) and (1.9) imply that $\gamma_{1}|_{\mathbb{S}^{d-1}}=\gamma_{2}|_{\mathbb{S}^{d-1}}$ and $\partial_{\nu}\gamma_{1}|_{\mathbb{S}^{d-1}}=\partial_{\nu}\gamma_{2}|_{\mathbb{S}^{d-1}}$. Lemma˜2.5 ensures then that $(\Delta\gamma^{\mathrm{B}}_{1})|_{U_{s}}=(\Delta\gamma^{\mathrm{B}}_{2})|_{U_{s}}$; therefore, by (1.5), $w:=\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2}$ solves $\Delta w=0$ on $U_{s}$ with $w|_{\mathbb{S}^{d-1}}=\partial_{\nu}w|_{\mathbb{S}^{d-1}}=0$. This implies $w|_{U_{s}}=0$; the result follows now from ˜2. ∎
Proof of ˜5.
By (2.18) we have that $\gamma^{\mathrm{B}}=a(\gamma)-2a(\gamma)T_{d}(V^{\mathrm{B}})$. Using this, the case ii) ($d=2$) and the regularity of $\gamma^{\mathrm{B}}_{\mathrm{reg}}$ in i) ($d\geq 3$) are immediate by Proposition˜2.2 and Lemma˜2.4 ii).
To finish the proof of i) notice that $T_{d}(|\cdot|^{-2\kappa-2})=\frac{r^{-2\kappa}-1}{2\kappa(d-2-2\kappa)}$ for all $0<\kappa<\nu_{d}$, and $T_{d}(|\cdot|^{-2})=-(d-2)^{-1}\log|x|$, $d\geq 3$. This yields using Proposition˜2.2 i) the representation in $d\geq 3$ with $c_{j}=-a_{j}\frac{1}{2\kappa(d-2-2\kappa_{j})}$ for $j=1,\dots,J$. Since $a_{j}\leq 0$ for all $j\geq 1$, then $c_{j}\geq 0$ for all $j\geq 1$. Finally iii) follows immediately from Proposition˜2.2 iii) using that $\gamma^{\mathrm{B}}$ and $\sqrt{\gamma}$ are, respectively, solutions of the elliptic equations $\Delta\gamma^{\mathrm{B}}=2a(\gamma)V^{\mathrm{B}}$ and $\Delta\sqrt{\gamma}=V\sqrt{\gamma}$. ∎
3. The Born approximation at the origin: singularity and resonances
3.1. Weyl-Titchmarsh functions and DtN maps
Write
$\nu_{d}:=\frac{d-2}{2},\qquad\mathbb{C}_{+}=\{z\in\mathbb{C}:\operatorname{\mathrm{Re}}(z)>0\}.$
In the presence of radial symmetry, it is possible to express $\Lambda^{\mathrm{p}}_{V}$ in terms of the Dirichlet-to-Neumann map of a one-dimensional Schrödinger operator on the half-line:
(3.1) $\begin{cases}-v_{z}^{\prime\prime}+Qv_{z}=-z^{2}v_{z},\qquad\text{ in }\mathbb{R}_{+},\\ v_{z}\in L^{2}(\mathbb{R}_{+}),\end{cases}$
When the potential $Q$ belongs to $L^{1}(\mathbb{R}_{+})$, it is known that there exists $\beta_{Q}>0$ such that, when $z\in\mathbb{C}_{+}\setminus[0,\beta_{Q}]$, problem (3.1) has a unique solution. One then defines the Weyl-Titchmarsh function of $Q$, or Dirichlet-to-Neumann map, as:
(3.2) $m_{Q}(-z^{2}):=\frac{v^{\prime}_{z}(0)}{v_{z}(0)}.$
When $Q\in L^{1}(\mathbb{R}_{+})$, it has been shown in Simon’s seminal paper [Sim99] that there exists a function $A_{Q}\in L^{1}_{\mathrm{loc}}(\mathbb{R}_{+})$, called the $A$-amplitude of $Q$, such that
$m_{Q}(-z^{2})=-z-\int_{0}^{\infty}e^{-2zt}A_{Q}(t)\,dt\qquad\text{for all }z\text{ such that }\operatorname{\mathrm{Re}}(z)>z_{Q},$
where $z_{Q}=\frac{\|Q\|_{1}}{2}>0$. As a matter of fact, it is proved in [MM24] that the $A$-amplitude coincides with the notion of the Born approximation for the inverse problem of recovering a potential from its Weyl-Titchmarsh function.
The connection of $m_{Q}$ with radial Dirichlet-to-Neumann maps is as follows. Given a potential $V\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$, with $p>d/2$ and $Y_{k}\in\mathfrak{H}^{d}_{k}$ then the unique solution $u\in H^{1}(\mathbb{B}^{d})$ of
(3.3) $\begin{cases}(-\Delta+V)u&=0,\quad\text{ in }\mathbb{B}^{d},\\ \hfill u|_{\mathbb{S}^{d-1}}&=Y_{k},\end{cases}$
can be written in the form
[…]
where $v_{k+\nu_{d}}\in L^{2}(\mathbb{R}_{+})$ solves the boundary value problem (3.1) for all $k\in\mathbb{N}_{0}$ with potential $Q=Q_{V}$ given by any of the following equivalent identities. If $V=q_{V}(|\cdot|)$ then
(3.5) $Q_{V}(t):=e^{-2t}q_{V}(e^{-t}),\;t\in\mathbb{R}_{+},\qquad\text{and}\qquad q_{V}(r)=r^{-2}Q_{V}(-\log r),\;r\in(0,1].$
Moreover
(3.6) $\lambda_{k}[V]=-m_{Q_{V}}\left(-(k+\nu_{d})^{2}\right)-\nu_{d}\qquad\forall k\in\mathbb{N}_{0}.$
In [DMMN24, Theorem 1] it was proved that there exists a $k_{V}>0$ such that
$\sigma_{k}[V^{\mathrm{B}}]=\lambda_{k}[V]-k,\qquad\forall k\in\mathbb{N}_{0},\;k>k_{V},$
where
(3.7) $V^{\mathrm{B}}(x):=|x|^{-2}A_{Q_{V}}(-\log|x|),\qquad x\in\mathbb{B}^{d}\setminus\{0\},$
This means that the $A$-amplitude of $Q_{V}$ determines the Born approximation of the potential $V$.
Remark 3.1.
The work of Remling [Rem03] implies that not all locally integrable functions can be $A$-amplitudes of Schrödinger operators. This implies, together with identities (2.19) and (3.7), that non-trivial necessary conditions are required to characterize the subset of $L^{1}_{\mathrm{rad}}(\mathbb{B}^{d};\mathbb{R})$ formed by the Born approximations $\gamma^{\mathrm{B}}$ corresponding to conductivities in $W^{2,p}_{\mathrm{rad}}(\mathbb{B}^{d})$. This means that ˜1 provides a partial, but not total, characterization of the set of rotation-invariant DtN maps.
The important formula (3.7) which connects the Born approximation $V^{\mathrm{B}}$ to the $A$-amplitude of the potential $Q_{V}$ will allow us to extract precise information on the singularities of $V^{\mathrm{B}}$ at 0 from certain asymptotic properties of $A_{Q_{V}}$.
3.2. Asymptotics of $A$-amplitudes
In this section, we review the main results from [RS00] that will be needed in the sequel.
Let $Q\in L^{1}(\mathbb{R}_{+},\mathbb{R})$ be such that
(3.8) $\int_{0}^{\infty}|Q(t)|(1+t)\,dt<\infty.$
We define the 11-$d$ Schrödinger operator $H_{Q}$, acting on $L^{2}(\mathbb{R}_{+})$ by
(3.9) $H_{Q}u(t):=\left(-\frac{d^{2}\,}{dt^{2}}+Q(t)\right)u(t),\qquad t\in\mathbb{R}_{+},$
with its domain (corresponding to Dirichlet boundary condition at 0)
$D(H_{Q})=\{u\in L^{2}(\mathbb{R}_{+})\,:\,u\in\mathcal{A}\mathcal{C}_{\mathrm{loc}}(\mathbb{R}_{+}),\,H_{Q}u\in L^{2}(\mathbb{R}_{+}),\,u(0)=0\}$
where $Q$ is determined by (3.5). Note that the operator $(H_{Q},D(H_{Q}))$ is self-adjoint on $L^{2}(\mathbb{R}_{+})$.
We now summarize well-known facts, see [RS00] and [Mar86]. Under the previous assumption, for every $z\in\overline{\mathbb{C}_{+}}$ there exists a unique solution $\psi(t,z)$ of (3.1) such that
(3.10) $\psi(t,z)=e^{-tz}(1+o(1)),\qquad t\to+\infty.$
These solutions are known as Jost solutions (they are usually defined using the variable $z^{\prime}=iz$), and $F(z):=\psi(0,z)$ is the Jost function associated to $Q$. Also, both functions $F$ and $\psi(t,\cdot)$ are analytic in $\mathbb{C}_{+}$. Notice that the zeros of $F$ at $\mathbb{C}_{+}$ occur at the points $z$ such that $z^{2}$ is a Dirichlet eigenvalue of $H_{Q}$.
One can show that $H_{Q}$ has no eigenvalues in $0,\infty)$, and that it always has a finite number of negative eigenvalues $\kappa_{j}^{2}$, where $j=1,\dots,J$, with $J\in\mathbb{N}_{0}$ and $0<\kappa_{J}<\dots<\kappa_{1}$. Recall that the Jost function associated to $Q$ satisfies $F(\kappa_{j})=0$ for all $j\in J$. It will also be relevant if the Jost function vanishes at $z=0$. In fact, as in [RS00], we will say that $H_{Q}$ has a zero resonance iff $F(0)=0$.
Let $\ell\geq 0$. Define the weighted $L^{1}_{\ell}(\mathbb{R}_{+})$ space with norm
(3.11) $\left\lVert f\right\rVert_{L^{1}_{\ell}(\mathbb{R}_{+})}:=\int_{0}^{\infty}|f(t)|(1+t)^{\ell}\,dt.$
Theorem 3.2 ([RS00, Theorem 1.3]).
Let $\ell\geq 3$, $Q\in L^{1}_{\ell}(\mathbb{R}_{+})$ and $(\kappa_{j})_{j=1}^{J}$ as above. Then, there exist some constants $(a_{j})_{j=0}^{J}$, and a function $g\in L^{1}_{\ell-1}(\mathbb{R}_{+})$ such that
(3.12) $A_{Q}(t)=a_{0}+\sum_{j=1}^{J}a_{j}e^{2t\kappa_{j}}+g(t).$
The constant $a_{0}$ vanishes if $H_{Q}$ has no zero resonances, and one always has $a_{1},\dots,a_{J}\leq 0$. As a consequence,
(3.13) $m_{Q}(-z^{2})=\int_{0}^{\infty}e^{-2zt}A_{Q}(t)\,dt,$
for all $z$ such that $\operatorname{\mathrm{Re}}(z)>\kappa_{1}$ if the point spectrum of $H_{Q}$ is non-empty, $\operatorname{\mathrm{Re}}(z)>0$ if it is empty, and $\operatorname{\mathrm{Re}}(z)\geq 0$ if, in addition, there are no zero resonances.
Proof.
The identity (3.12) is a direct consequence of [RS00, Theorem 1.3]. The constants satisfy $a_{1},\dots,a_{J}\leq 0$ as follows from [RS00, Remark 3, p. 321]. The fact that (3.13) holds for all $\operatorname{\mathrm{Re}}(z)>\kappa_{1}$ follows by the analyticity of both sides of the identity in $\{z\in\mathbb{C}:\operatorname{\mathrm{Re}}(z)>\kappa_{1}\}$. In the case of empty point spectrum, the same holds in $\{z\in\mathbb{C}:\operatorname{\mathrm{Re}}(z)>0\}$. Further, if there are no zero resonances, then $A_{Q}(t)$ belongs to $L^{1}(\mathbb{R}_{+})$ and both sides of (3.13) are continuous in $\{z\in\mathbb{C}:\operatorname{\mathrm{Re}}(z)\geq 0\}$, so they must be equal. ∎
3.3. Singularities of the Born approximation
We will now describe how the results of the previous section translate into results on the Born approximation for Schrödinger operator Dirichlet-to-Neumann maps. The following Lemma shows, in particular, that as soon as $p>d/2$ and $V\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$ we have that $Q_{V}$ satisfies (3.8).
Lemma 3.3.
Let $p>d/2$ and $\alpha\geq 0$. Then, there exists a constant $C(\alpha,p,d)>0$ such that, for every $V\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d})$, the potential $Q_{V}$ given by (3.5) satisfies
$\left\lVert Q_{V}\right\rVert_{L^{1}_{\alpha}(\mathbb{R}_{+})}\leq C(\alpha,p,d)\left\lVert V\right\rVert_{L^{p}(\mathbb{B}^{d})}.$
In particular, $C(0,p,d)=\left(\frac{p-1}{2p-d}\right)^{1-1/p}$.
Proof.
Let $1/p+1/q=1$. Then
$\left\lVert Q_{V}\right\rVert_{L^{1}_{\alpha}(\mathbb{R}_{+})}=\int_{\mathbb{R}_{+}}|Q_{V}(t)|(1+t)^{\alpha}\,dt$ $=\int_{\mathbb{B}^{d}}|V(x)|(1+|\log|x||)^{\alpha}|x|^{2-d}\,dx$
$\leq\left\lVert V\right\rVert_{L^{p}(\mathbb{B}^{d})}\left(\int_{\mathbb{B}^{d}}(1+|\log|x||)^{q\alpha}|x|^{q(2-d)}\,dx\right)^{1/q},$
and the second integral is finite for all $\alpha\geq 0$ and $1\leq q<d/(d-2)$, which implies $p>d/2$. Computing explicitly the integral when $\alpha=0$ yields the first inequality. ∎
As a consequence of this result and Theorem˜3.2 we obtain the following description of the singularity at the origin of the Born approximation. Recall from previous sections that
(3.14) $\left\lVert f\right\rVert_{d,\alpha}:=\int_{\mathbb{B}^{d}}|f(x)||x|^{2-d}(1-\log|x|)^{\alpha}\,dx<\infty,\qquad\alpha\geq 0,\,d\geq 2.$
Theorem 3.4.
Let $d\geq 2$, $V\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$ with $p>d/2$, and $(\kappa_{j})_{j=1}^{J}$ such that $(-\kappa_{j}^{2})_{j=1}^{J}$ is the point spectrum of $H_{Q_{V}}$, as above. There exist some constants $a_{0}\in\mathbb{R},\,a_{1},\dots,a_{J}\leq 0$, such that
$V^{\mathrm{B}}(x)=\frac{a_{0}}{|x|^{2}}+\sum_{j=1}^{J}\frac{a_{j}}{|x|^{2\kappa_{j}+2}}+V^{\mathrm{B}}_{\mathrm{reg}}(x),$
where $V^{\mathrm{B}}_{\mathrm{reg}}$ is a radial function satisfying $\left\lVert V^{\mathrm{B}}_{\mathrm{reg}}\right\rVert_{d,\alpha}<\infty$ for all $\alpha>0$.
In particular, $a_{0}=0$ if $H_{Q_{V}}$ has no zero resonance.
Proof.
Let $Q_{V}$ be given by (3.9). Then $Q\in L^{1}_{\ell}(\mathbb{R}_{+})$ for all $\ell\geq 1$, by Lemma˜3.3. Therefore it satisfies the conditions to apply Theorem˜3.2. Thus
$A_{Q_{V}}(t)=a_{0}+\sum_{j=1}^{J}a_{j}e^{2t\kappa_{j}}+g(t),$
for some constants $(a_{j})_{j=0}^{J}$. Here $g\in L^{1}_{\ell}(\mathbb{R}_{+})$ for all $\ell\geq 0$. The constant $a_{0}=0$ if $H_{Q_{V}}$ has no zero resonances. Thus, the statement for $V^{\mathrm{B}}$ follows directly using (3.7), with $V^{\mathrm{B}}_{\mathrm{reg}}(x):=|x|^{-2}g(-\log|x|)$. ∎
In general, each constant $a_{j}$ is explicitly related to certain values of the eigenfunction with eigenvalue $\kappa_{j}$, see [RS00].
Theorem˜3.4 has interesting consequences. In the first place, it shows that $V^{\mathrm{B}}$ can be a very singular object: one can always find smooth radial potentials $V$ such that $H_{Q_{V}}$ has an eigenvalue $\kappa_{1}$ as large as desired (for example, a smooth potential with large negative values). On the other hand, if one imposes conditions on the potentials that bound the value of $\kappa_{1}$, then the Born approximation can be an integrable function. We will examine this in the following section.
Remark 3.5.
If $V\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$, $p>d/2$ and $V(x)\geq 0$ $a.e.$ on $\mathbb{B}^{d}$, then $H_{Q_{V}}$ has no bound states nor zero resonances. As a consequence of Theorem˜3.4 we see that $\left\lVert V^{\mathrm{B}}\right\rVert_{d,\alpha}<\infty$ for all $\alpha\geq 0$, which is stronger that $V^{\mathrm{B}}\in L^{1}(\mathbb{B}^{d})$.
Remark 3.6.
Theorem˜3.4 can be improved to take into account the complex resonances of $H_{Q_{V}}$, which will add oscillatory terms to the asymptotic expansion. If one assumes that for some $\infty<\delta<0$ the potential satisfies $\int_{\mathbb{B}^{d}}|V(x)||x|^{2-d+\delta}\,dx<\infty,$ (which always holds if $V\in L^{p}(\mathbb{B}^{d})$ with $p>d/2$) then, for a fixed $\delta<\delta^{\prime}<\infty$,
$V^{\mathrm{B}}(x)=\sum_{j=1}^{J}\frac{a_{j}}{|x|^{2\kappa_{j}+2}}+\sum_{j=1}^{M}\frac{b_{j}}{|x|^{2\lambda_{j}+2}}+\sum_{j=1}^{N}c_{j}\frac{\cos(\theta_{j}-2\nu_{j}\log|x|)}{|x|^{2\mu_{j}+2}}+V^{\mathrm{B}}_{\mathrm{reg}}(x),$
where $0<\kappa_{J}<\dots<\kappa_{1}$ corresponds to the negative eigenvalues $\kappa_{j}^{2}$ of $H_{Q_{V}}$, $\delta^{\prime}<\lambda_{M}\leq\dots\leq\lambda_{1}\leq 0$ corresponds to the real resonances larger than $\delta^{\prime}$, $\mu_{j}\pm i\nu_{j}$ with $\delta^{\prime}<\mu_{j}<0$ correspond for all $j=1,\dots,N$ to the complex resonances with real part larger than $\delta^{\prime}$, and, finally, $V^{\mathrm{B}}_{\mathrm{reg}}(x)$ is a radial function satisfying that $\int_{\mathbb{B}^{d}}|x|^{2-d+\delta^{\prime}}|V^{\mathrm{B}}_{\mathrm{reg}}(x)|\,dx<\infty$. This follows from [RS00, Theorem 1.4] using the same strategy as in the proof of Theorem˜3.4. This oscillatory behavior has been observed in the numerical reconstructions of $V^{\mathrm{B}}(x)$ in [BCMM24].
Another consequence of Theorem˜3.2 is the following.
Lemma 3.7.
Let $d\geq 2$, $V\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$ with $p>d/2$. Then
(3.15) $\int_{\mathbb{B}^{d}}|V^{\mathrm{B}}(x)||x|^{2k}\,dx<\infty,\quad\text{and}\quad\sigma_{k}[V^{\mathrm{B}}]=\lambda^{\mathrm{p}}_{k}[V]-k,\qquad\forall k\in\mathcal{N},$
where $\mathcal{N}\subseteq\mathbb{N}_{0}$ is given by:
  1. i)
    If the point spectrum of $H_{Q_{V}}$ is non-empty then $\mathcal{N}=\{k\in\mathbb{N}_{0}:k>\kappa_{1}-\nu_{d}\}$.
  2. ii)
    If the point spectrum of $H_{Q_{V}}$ is empty then:
    1. (a)
      If $d>2$ then $\mathcal{N}=\mathbb{N}_{0}$.
    2. (b)
      If $d=2$ then $\mathcal{N}=\mathbb{N}_{0}$ provided that $H_{Q_{V}}$ has no zero resonance, and $\mathcal{N}=\mathbb{N}$ otherwise.
Proof.
Let $Q_{V}$ be given by (3.9). Then $Q_{V}\in L^{1}_{\ell}(\mathbb{R})$ for all $\ell\geq 3$ by Lemma˜3.3, so Theorem˜3.2 holds. Then, by (3.6), (3.13) and (3.7) in that order:
$\lambda_{k}[V]-k=-m_{Q_{V}}\left(-(k+\nu_{d})^{2}\right)-(k+\nu_{d})\\ =\int_{0}^{\infty}e^{-2kt}A_{Q_{V}}(t)e^{-(d-2)t}\,dt=\frac{1}{|\mathbb{S}^{d-1}|}\int_{\mathbb{B}^{d}}V^{\mathrm{B}}(x)|x|^{2k}\,dx=\sigma_{k}[V^{\mathrm{B}}],$
where the integrals are always absolutely convergent. Notice that, by Theorem˜3.2, the second equality holds when $k+\nu_{d}>\kappa_{1}$ if the spectrum of $H_{Q_{V}}$ is non-empty, when $k+\nu_{d}>0$ if it is empty, and when $k+\nu_{d}\geq 0$ if, in addition, there are no zero resonances. This yields all the cases enumerated in the statement. ∎
3.4. The conductivity case
If the potential $V$ arises from a conductivity, then the operator $H_{Q_{V}}$ is always uniformly bounded from below and $\kappa_{1}<\nu_{d}$, as we now show. As a consequence, the singularities of the Born approximation will be of integrable type in this case. We first need the following lemma.
Lemma 3.8.
Assume $Q_{V}\in{L}^{1}_{2}(\mathbb{R}_{+})$. Then 0 is not an eigenvalue of $H_{Q_{V}}$.
Proof.
At zero energy, any solution $u$ to $H_{Q_{V}}u=0$ on $\mathbb{R}_{+}$ satisfies the Volterra equation
$u(x)=A+Bx+\int_{x}^{\infty}(t-x)\,Q_{V}(t)\,u(t)\,dt.$
Writing w(x)=u(x)(A+Bx)w(x)=u(x)-(A+Bx), we obtain, using that $t-x\leq t$ and $(1+t^{2})Q_{V}\in L^{1}(\mathbb{R}_{+})$,
$|w(x)|\leq(|A|+|B|)\int_{x}^{\infty}t(1+t)|Q_{V}(t)|\,dt+\!\int_{x}^{\infty}t|Q_{V}(t)|\,|w(t)|\,dt.$
Gronwall’s lemma gives:
$|w(x)|\leq(|A|+|B|)M(x)\exp\left(\int_{x}^{\infty}t|Q_{V}(s)|ds\right),\qquad M(x):=\int_{x}^{\infty}t(1+t)|Q_{V}(t)|\,dt,$
since $M(x)$ is a decreasing function. This shows that $w(x)\to 0$ as $x\to\infty$. If $u\in L^{2}(\mathbb{R}_{+})$ this forces $A=B=0$, and therefore $w=0$. ∎
Lemma 3.9.
Let $p>d/2$ and $\gamma\in W^{2,p}_{\mathrm{rad}}(\mathbb{B}^{d})$ be a conductivity, and let $V\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d})$ be given by (2.5). Let $H_{Q_{V}}$ be the associated Hamiltonian defined in (3.9). Then
  1. i)
    If $d>2$, the discrete spectrum of $H_{Q_{V}}$ is contained in $(-\nu_{d}^{2},0)$.
  2. ii)
    If $d=2$, the discrete spectrum of $H_{Q_{V}}$ is empty and it has no zero resonance.
Proof.
Let $\gamma(x)$ such that $\gamma(x)=\gamma_{0}(|x|)$, and define $\phi(t):=e^{-\nu_{d}t}\sqrt{\gamma_{0}(e^{-t})}$. From (2.5), (3.9), (3.5) and straightforward computations, it follows that
$Q_{V}(t)=\frac{\phi^{\prime\prime}(t)}{\phi(t)}-\nu_{d}^{2}.$
Let $w(t):=\frac{\phi^{\prime}(t)}{\phi(t)}=(\log\phi(t))^{\prime}$. Then $Q_{V}$ can be written explicitly in terms of $w$ as
$Q_{V}(t)=\left(\frac{\phi^{\prime}(t)}{\phi(t)}\right)^{\prime}+\left(\frac{\phi^{\prime}(t)}{\phi(t)}\right)^{2}-\nu_{d}^{2}=w^{\prime}(t)+w(t)^{2}-\nu_{d}^{2}.$
With this representation for $Q_{V}$, one can prove that one has the factorization
$H_{Q_{V}}+\nu_{d}^{2}=B^{*}B,\qquad B:=-\partial_{t}+w(t),\quad B^{*}=\partial_{t}+w(t).$
Consequently, for every $u\in H^{1}_{0}(\mathbb{R}_{+})$, one has that
$(u,(H_{Q_{V}}+\nu_{d}^{2})u)_{L^{2}(\mathbb{R}_{+})}=\int_{0}^{\infty}|-\partial_{t}u(t)+w(t)u(t)|^{2}\,dt=\int_{0}^{\infty}\phi(t)^{2}\left|\left(\frac{u(t)}{\phi(t)}\right)^{\prime}\right|^{2}\,dt\geq 0.$
Notice that $(u,(H_{Q_{V}}+\nu_{d}^{2})u)_{L^{2}(\mathbb{R}_{+})}$ is always positive in $H^{1}_{0}(\mathbb{R}_{+})$ since the second integral only vanishes for $u=c\phi$ with $c\in\mathbb{C}$, but $\phi(0)=\gamma_{0}(1)>0$, hence $\phi\notin H^{1}_{0}(\mathbb{R}_{+})$. Therefore
$\operatorname{\mathrm{Spec}}_{H^{1}_{0}}(H_{Q_{V}})\subseteq(-\nu_{d}^{2},\infty).$
By Lemma˜3.8, 0 is not an eigenvalue $H_{Q_{V}}$. By [Mar86, Lemma 3.1.1], for $\kappa\in\mathbb{R}$ the solutions of $v^{\prime\prime}+Q_{V}v=\kappa^{2}v_{z}$ has two independent solutions $v_{\pm}$ that behave as $e^{\pm i\kappa t}$ for $t\to+\infty$, so $\kappa^{2}\in(0,\infty)$ can never be an eigenvalue of $H_{Q_{V}}$. This ensures the absence of embedded eigenvalues in $0,\infty)$ (since $Q_{V}\in L^{1}_{\alpha}(\mathbb{R}_{+})\subset L^{1}(\mathbb{R}_{+})$, one can also apply the classical result of Kneser–Weyl [Kne26, Wey10] –see also Simon [Sim19]– to prove the absence of embedded eigenvalues in $(0,\infty)$). Hence, it follows that the discrete spectrum of $H_{Q_{V}}$ must be contained in $(-\nu_{d}^{2},0)$ if $d>2$, and must be empty if $d=2$.
It remains to prove that there is not a zero resonance when $d=2$. Notice that
$(H_{Q_{V}}+\nu_{d}^{2})\phi=0,\qquad d\geq 2,$
so if $d=2$ we have $H_{Q_{V}}\phi=0$. Now, $\phi(t)\to\gamma(0)>0$ as $t\to\infty$, so $\phi(t)=\psi(t,0)$ is the Jost solution with $z=0$ associated to $H_{Q_{V}}$ introduced in Section˜3.2. The Jost solution is unique and we have $\psi(0,0)=\phi(0)=\gamma_{0}(1)>0$. This implies that $H_{Q_{V}}$ has no Dirichlet zero resonance. ∎
As the following proposition shows, the previous conditions on the spectrum are enough to prove that $V^{\mathrm{B}}\in L^{1}(\mathbb{B}^{d})$ always, with some extra regularity properties.
Proposition 3.10.
Let $V\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$ with $p>d/2$. Let $H_{Q_{V}}$ be the operator defined in (3.9) and $V^{\mathrm{B}}$ as in (3.7). The following assertions hold:
  1. i)
    $V^{\mathrm{B}}\in L^{1}_{\mathrm{rad}}(\mathbb{B}^{d})$ and $\sigma_{k}[V^{\mathrm{B}}]=\lambda^{\mathrm{p}}_{k}[V]-k$, for all $k\in\mathbb{N}_{0}$.
  2. ii)
    If $d>2$ and the discrete spectrum $(-\kappa_{j}^{2})_{j=1}^{J}$ of $H_{Q_{V}}$ is contained in $\left(-\nu_{d}^{2},0\right)$, then there exist some constants $a_{0}\in\mathbb{R},\,a_{1},\dots,a_{J}\leq 0$, such that
    $V^{\mathrm{B}}(x)=\frac{a_{0}}{|x|^{2}}+\sum_{j=1}^{J}\frac{a_{j}}{|x|^{2\kappa_{j}+2}}+V^{\mathrm{B}}_{\mathrm{reg}}(x),$
    where $\left\lVert V^{\mathrm{B}}_{\mathrm{reg}}\right\rVert_{d,\alpha}<\infty$ for all $\alpha\geq 0$ (see (3.14) for the definition of $\left\lVert\cdot\right\rVert_{d,\alpha}$).
  3. iii)
    $d=2$, the discrete spectrum of $H_{Q_{V}}$ is empty and it has no zero resonance. Then $V^{\mathrm{B}}$ satisfies $\left\lVert V^{\mathrm{B}}\right\rVert_{2,\alpha}<\infty$ for all $\alpha\geq 0$.
Proof.
Statement i) follows from Lemma˜3.7. ii) and iii) follow directly from Theorem˜3.4 and Lemma˜3.9 (for $H_{Q_{V}}$ the discrete spectrum and the point spectrum coincide, since the essential spectrum is $0,\infty)$ and 0 is never an eigenvalue). ∎
3.5. Explicit examples
We conclude this section by presenting a family of conductivities for which the Born approximation can be computed explicitly.
Let $d\geq 2$, $\mu>0$ and $\nu\geq 0$. Define
$\rho_{d,\mu,\nu}(r):=\begin{cases}\displaystyle\frac{1}{\nu+\nu_{d}}\left(\frac{2\mu}{1+\alpha r^{2\mu}}+\nu_{d}-\mu\right),&(d,\nu)\neq(2,0)\vskip 6.0pt plus 2.0pt minus 2.0pt\\ \displaystyle\mu\frac{1-r^{2\mu}}{1+r^{2\mu}},&(d,\nu)=(2,0)\end{cases},\qquad\alpha:=\frac{\mu-\nu}{\mu+\nu},$
and note that $\rho_{d,\mu,\nu}(r)>0$ for every $r\in(0,1)$.
Proposition 3.11.
Let $d\geq 2$, $\mu>0$ and $\nu\geq 0$ ($\nu>0$ when $d=2$) and set
(3.16) $\gamma_{d,\mu,\nu}(x):=\rho_{d,\mu,\nu}(|x|)^{2}.$
Then $\gamma_{d,\mu,\nu}$ is a conductivity in $W^{2,p}(\mathbb{B}^{d})$ for some $p>d/2$ (every $p$ when $\mu\geq 1$) such that $\gamma_{d,\mu,\nu}|_{\mathbb{S}^{d-1}}=1$ and whose Born approximation is given by
(3.17) $\gamma^{\mathrm{B}}_{d,\mu,\nu}(x)=\begin{cases}\displaystyle 1+\frac{(\nu^{2}-\mu^{2})}{\nu(\nu+\nu_{d})}(|x|^{2\nu}-1),&\nu>0,\\[10.0pt] \displaystyle 1-\frac{2\mu^{2}}{\nu_{d}}\,\log|x|,&\nu=0,\;d\geq 3.\end{cases}$
Proof.
Direct computation shows that:
$\Delta\sqrt{\gamma_{d,\mu,\nu}}=V\sqrt{\gamma_{d,\mu,\nu}},\qquad V(x):=-8\mu^{2}\alpha\frac{|x|^{2(\mu-1)}}{\left(1+\alpha|x|^{2\mu}\right)^{2}}.$
In [DMMN24], it is shown that the Born approximation of $V$ is given by
$V^{\mathrm{B}}(x)=2(\nu^{2}-\mu^{2})|x|^{2(\nu-1)}.$
These potentials correspond, after performing the change of variables (3.5), to those studied in [GS00, Section 11]. One can also check, by direct computation, that if $\gamma^{\mathrm{B}}$ is given by (3.17) then
$\Delta\gamma^{\mathrm{B}}_{d,\mu,\nu}=2V^{\mathrm{B}},\qquad\gamma^{\mathrm{B}}_{d,\mu,\nu}|_{\mathbb{S}^{d-1}}=1.$
The claim then follows from relation (2.19) and uniqueness of solutions to the Poisson equation. ∎
This example illustrates some of the results in this section.
Remark 3.12.
  1. i)
    The parameter $\nu$ represents the unique real resonance of $H_{Q_{V}}$, (see [GS00], case 2, p. 636).
  2. ii)
    The presence of the zero resonance for $d\geq 3$ results in a Born approximation $\gamma^{\mathrm{B}}$ with a logarithmic singularity at the origin (and $V^{\mathrm{B}}$ with a $|x|^{-2}$ singularity), as predicted by Proposition˜3.10 and ˜5.
  3. iii)
    In the case $d=2$ and $\nu=0$ there is a zero resonance and $\gamma_{2,\mu,0}$ is a degenerate conductivity, since it vanishes on $\mathbb{S}^{1}$. This is consistent with Lemma˜3.9 ii).
4. Stability properties of the Born approximation
4.1. Hölder stability
In the Calderón problem one can prove estimates of the kind
(4.1) $\left\lVert\gamma_{1}-\gamma_{2}\right\rVert_{L^{q}(\mathbb{B}^{d})}\leq\omega_{\mathcal{K}}\left(\left\lVert\Lambda_{\gamma_{1}}-\Lambda_{\gamma_{2}}\right\rVert_{\mathcal{L}(H^{1/2},H^{-1/2})}\right),\qquad\gamma_{1},\gamma_{2}\in\mathcal{K},$
where $\mathcal{K}$ is a compact subset of $L^{q}(\mathbb{B}^{d})$, $1\leq q\leq\infty$, and $\omega_{\mathcal{K}}$ is a modulus of continuity [Ale88, BBR01, CFR10, San13]. Assuming the set $\mathcal{K}$ only contains radial functions, this becomes
$\left\lVert\gamma_{1}-\gamma_{2}\right\rVert_{L^{q}(\mathbb{B}^{d})}\leq\omega_{\mathcal{K}}\left(\sup_{k\in\mathbb{N}}|\lambda_{k}[\gamma_{1}]-\lambda_{k}[\gamma_{2}]|\right),\qquad\gamma_{1},\gamma_{2}\in\mathcal{K}.$
Due to the strong continuity properties of the moment problem, the previous estimate implies
(4.2) $\left\lVert\gamma_{1}-\gamma_{2}\right\rVert_{L^{q}(\mathbb{B}^{d})}\leq\omega_{\mathcal{K}}\left(\left\lVert\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2}\right\rVert_{W^{2,1}(\mathbb{B}^{d})}\right),$
provided that $\gamma_{1},\gamma_{2}\in\mathcal{K}\cap W^{2,p}(\mathbb{B}^{d})$, for some $p>d/2$, and that $\gamma_{1}|_{\mathbb{S}_{d-1}}=\gamma_{2}|_{\mathbb{S}_{d-1}}$. Indeed, by ˜1 and (2.11) we have
$|\lambda_{k}(\gamma_{1})-\lambda_{k}$ $(\gamma_{2})|\leq\left|\sigma_{k}[\Delta(\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2})]\right|+\left|\partial_{\nu}\gamma_{1}|_{\mathbb{S}_{d-1}}-\partial_{\nu}\gamma_{2}|_{\mathbb{S}_{d-1}}\right|$
$=\left|\sigma_{k}[\Delta(\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2})]\right|+\left|\sigma_{0}[\Delta(\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2})]\right|$
$\leq 2\sup_{k\in\mathbb{N}}\left|\sigma_{k}[\Delta(\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2})]\right|\leq 2\left\lVert\Delta(\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2})\right\rVert_{L^{1}(\mathbb{B}^{d})}\leq 2\left\lVert\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2}\right\rVert_{W^{2,1}(\mathbb{B}^{d})}.$
It is well known that in general $\omega_{\mathcal{K}}$ in (4.1) cannot be taken better than logarithmic when $\mathcal{K}$ includes a ball in a Sobolev space, [Man01, KRS21]. ˜4 improves estimate (4.2) to a Hölder modulus of continuity, and gets rid of the assumption $\gamma_{1}|_{\mathbb{S}_{d-1}}=\gamma_{2}|_{\mathbb{S}_{d-1}}$.
4.2. Proof of ˜4
To get the estimate for the conductivity case, we use the following lemma, and two stability results for the Schrödinger problem from [DMMN24], which in turn are based on stability properties of the $A$-amplitude [DMMN24, Theorem 5.1]. Let $0<s<1$ and recall the definition of $U_{s}$ in (1.8).
Lemma 4.1.
Let $K>1$, $N>0$, $d/2<p<\infty$, $0<s<1$. There exist some constants $C_{K,N,p,d,s},C_{K,N,p,d}>0$ such that, for all $\gamma_{j}\in W^{2,p}_{\mathrm{rad}}(\mathbb{B}^{d})$, satisfying
(4.3) $K^{-1}\leq\gamma_{j}(x)\leq K,\qquad\left\lVert\gamma_{j}\right\rVert_{W^{2,p}(\mathbb{B}^{d})}\leq N,\quad j=1,2,$
then, if $a_{j}:=\gamma_{j}|_{\mathbb{S}^{d-1}}$, $b_{j}:=\partial_{\nu}\gamma_{j}|_{\mathbb{S}^{d-1}}$ one has the local estimate
$C_{K,N,p,d,s}^{-1}\left\lVert\gamma_{1}-\gamma_{2}\right\rVert_{W^{2,1}(U_{s})}\leq\left\lVert V_{1}-V_{2}\right\rVert_{L^{1}(U_{s})}+|a_{1}-a_{2}|+|b_{1}-b_{2}|.$
and the global estimate
$C_{K,N,p,d}^{-1}\left(\left\lVert\gamma_{1}-\gamma_{2}\right\rVert_{W^{2,1}(\mathbb{B}^{d})}+\left\lVert\gamma_{1}-\gamma_{2}\right\rVert_{H^{1}(\mathbb{B}^{d})}\right)\leq\left\lVert V_{1}-V_{2}\right\rVert_{L^{1}(\mathbb{B}^{d})}^{\beta}+|a_{1}-a_{2}|,$
for $\beta=\frac{p-r_{d}}{r_{d}(p-1)}$, with $r_{d}:=\frac{2d}{d+2}$ for $d\geq 3$, $r_{2}:=(p+1)/2$.
The proof of the lemma is rather technical but standard. For the sake of completeness, we include it in Appendix˜A.
Theorem 4.2 (Theorem 3 of [DMMN24]).
Let $d\geq 2$, $d/2<p\leq\infty$, and $0<s<1$. For all $N>|\mathbb{S}^{d-1}|^{1/p}$, there exist constants $C_{d,N,p,s}>0$ and $0<\delta_{d,p,s}<1$ such that, for all $V_{1},V_{2}\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d};\mathbb{R})$ satisfying
(4.4) $\max_{j=1,2}\left\lVert V_{j}\right\rVert_{L^{p}(U_{s})}\leq N,\qquad\qquad\left\lVert V^{\mathrm{B}}_{1}-V^{\mathrm{B}}_{2}\right\rVert_{L^{1}(U_{s})}<\delta_{d,p,s},$
one has
(4.5) $\left\lVert V_{1}-V_{2}\right\rVert_{L^{1}(U_{s})}<C_{d,N,p,s}\,\left\lVert V^{\mathrm{B}}_{1}-V^{\mathrm{B}}_{2}\right\rVert_{L^{1}(U_{s})}^{\frac{p}{2p-1}}.$
Theorem 4.3 (Theorem 4 of [DMMN24]).
Let $d\geq 2$, $d/2<p<\infty$ and M>d1M>d-1. There exists $C_{d,M,p}>0$ and $0<\alpha<1$, depending only on $M,d,p$, such that for every $V_{1},V_{2}\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d};\mathbb{R})$ satisfying
(4.6) $\max_{j=1,2}\left\lVert V_{j}\right\rVert_{L^{p}(\mathbb{B}^{d})}<\frac{|\mathbb{S}^{d-1}|^{\frac{1}{p}}}{4}\frac{p-d/2}{p-1}M,$
$\int_{\mathbb{B}^{d}}|V^{\mathrm{B}}_{1}(x)-V^{\mathrm{B}}_{2}(x)||x|^{M-(d-2)}\,dx<|\mathbb{S}^{d-1}|,$
one has
(4.7) $\left\lVert V_{1}-V_{2}\right\rVert_{L^{1}(\mathbb{B}^{d})}<C_{d,M,p}\left(\int_{\mathbb{B}^{d}}|V^{\mathrm{B}}_{1}(x)-V^{\mathrm{B}}_{2}(x)||x|^{M-(d-2)}\,dx\right)^{\alpha}.$
Lemma 4.4 (Lemma 5.6 of [DMMN24]).
Let $d\geq 2$, $N>0$ and $d/2<p<\infty$. There exists a constant $\beta(d,p)$ such that if $\left\lVert V\right\rVert_{L^{p}(\mathbb{B}^{d})}<N$, then $\left\lVert V^{\mathrm{B}}\right\rVert\leq\beta(d,p)N$.
Proof of ˜4.
Since the conductivities are radial and (1.5) holds, by Lemma˜2.3 ii) we have for some $C>0$ that
(4.8) $|a_{1}-a_{2}|+|b_{1}-b_{2}|\leq C\left\lVert\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2}\right\rVert_{W^{2,1}(\mathbb{B}^{d})},$
where $a_{j}:=\gamma_{j}|_{\mathbb{S}^{d-1}}$, $b_{j}:=\partial_{\nu}\gamma_{j}|_{\mathbb{S}^{d-1}}$. Locally, the same bound holds:
(4.9) $|a_{1}-a_{2}|+|b_{1}-b_{2}|\leq\frac{C(d,s)}{1-s}\left\lVert\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2}\right\rVert_{W^{2,1}(U_{s})}.$
Using the previous estimates, the theorem follows combining Lemma˜4.1 with Theorems 4.2 and 4.3 as we now show.
Conditions (1.10) on $\gamma_{j}$, $j=1,2$ imply a corresponding bound for $\left\lVert V_{j}\right\rVert_{L^{p}}$ using that $\Delta\gamma_{j}=V_{j}\sqrt{\gamma_{j}}$, so the first estimates in (4.4) and (4.6) hold. We now consider the global and local cases separately.
Proof of local estimate i). First we claim that for any given $0<\varepsilon<1$, there exists a $0<\delta<1$ (depending on $\varepsilon,d,K,N,s$) such that $\left\lVert V^{\mathrm{B}}_{1}-V^{\mathrm{B}}_{2}\right\rVert_{L^{1}(U_{s})}<\varepsilon$ holds if $\left\lVert\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2}\right\rVert_{W^{2,1}(\mathbb{B}^{d})}<\delta$ holds for $\delta$. This can be proved using (2.19) together with $K^{-1}\leq\gamma_{j}(x)\leq K$ which imply that
$\left\lVert V^{\mathrm{B}}_{1}-V^{\mathrm{B}}_{2}\right\rVert_{L^{1}(U_{s})}=\left\lVert(2a_{1})^{-1}(2a_{1}V^{\mathrm{B}}_{1}-2a_{2}V^{\mathrm{B}}_{2})+a_{2}V^{\mathrm{B}}_{2}(a_{1}^{-1}-a_{2}^{-1})\right\rVert_{L^{1}(U_{s})}\\ \leq\frac{K}{2}\left\lVert\Delta\gamma^{\mathrm{B}}_{1}-\Delta\gamma^{\mathrm{B}}_{2}\right\rVert_{L^{1}(U_{s})}+|a_{1}^{-1}-a_{2}^{-1}||a_{2}|\left\lVert V^{\mathrm{B}}_{2}\right\rVert_{L^{1}(U_{s})},\$
we now use that $|a_{1}^{-1}-a_{2}^{-1}|\leq C_{K}|a_{1}-a_{2}|$, since $K^{-1}\leq a_{j}\leq K$, Lemma˜4.4 to bound the norm of $V^{\mathrm{B}}_{2}$, and (4.9), to obtain
(4.10) $\left\lVert V^{\mathrm{B}}_{1}-V^{\mathrm{B}}_{2}\right\rVert_{L^{1}(U_{s})}\leq C(d,K,N,s)\left\lVert\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2}\right\rVert_{W^{2,1}(U_{s})},$
which proves the claim. This proves that (4.4) holds, so that the conditions to apply Theorem˜4.2 are met. Inserting (4.10) and the local estimate in Lemma˜4.1 in (4.5) yields
$\left\lVert\gamma_{1}-\gamma_{2}\right\rVert_{W^{2,1}(U_{s})}<C_{N,K,d,p,s}\,\left\lVert\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2}\right\rVert_{W^{2,1}(U_{s})}^{\frac{p}{2p-1}}+|a_{1}-a_{2}|+|b_{1}-b_{2}|.$
Using (4.9) finishes the proof of the estimate.
Proof of global estimate ii) Using (1.10) and (4.8), the same arguments of the local case can be used to prove that
(4.11) $\left\lVert V^{\mathrm{B}}_{1}-V^{\mathrm{B}}_{2}\right\rVert_{L^{1}(\mathbb{B}^{d})}\leq C(d,K,N)\left\lVert\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2}\right\rVert_{W^{2,1}(\mathbb{B}^{d})}.$
Therefore, there exits a $0<\delta<1$ such that (4.7) holds if $\left\lVert\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2}\right\rVert_{W^{2,1}(\mathbb{B}^{d})}<\delta$ holds (the weight $|x|^{M-(d-2)}$ can be discarded since M>d2M>d-2). This proves that the conditions to apply Theorem˜4.3 are met. Inserting (4.10) and the global estimate of Lemma˜4.1 in (4.5) yields
$\left\lVert\gamma_{1}-\gamma_{2}\right\rVert_{H^{1}(\mathbb{B}^{d})}+\left\lVert\gamma_{1}-\gamma_{2}\right\rVert_{W^{2,1}(\mathbb{B}^{d})}<C_{N,K,d,p}\,\left\lVert\gamma^{\mathrm{B}}_{1}-\gamma^{\mathrm{B}}_{2}\right\rVert_{W^{2,1}(\mathbb{B}^{d})}^{\alpha\beta}+|a_{1}-a_{2}|,$
since the weight $|x|^{M-(d-2)}$ can be removed from Lemma˜4.1 using M>d2M>d-2 and that $\gamma^{\mathrm{B}}\in L^{1}(\mathbb{B}^{d})$. The Hölder exponent of this estimate is just the product of the one of Theorem˜4.3, with the one of Lemma˜4.1. Finally, one can use (4.8) to bound $|a_{1}-a_{2}|$, and that, for all $1<q<p$,
$\left\lVert\gamma_{1}-\gamma_{2}\right\rVert_{W^{2,q}(\mathbb{B}^{d})}\leq C_{N,d,p,q}\left\lVert\gamma_{1}-\gamma_{2}\right\rVert_{W^{2,1}(\mathbb{B}^{d})}^{\frac{p-q}{q(p-1)}},$
by interpolation, since $\left\lVert\gamma_{j}\right\rVert_{W^{2,p}(\mathbb{B}^{d})}\leq N$, with $j=1,2$. This finishes the proof of the Theorem. ∎
Appendix A Proof of Lemma˜4.1
Global estimate: In this proof, $C$ will always denote a constant that depends on $d,p,K,N$.
Let $v=\sqrt{\gamma_{1}}-\sqrt{\gamma_{2}}$. Then, by (2.5) one has that $\Delta v+V_{1}v=-\sqrt{\gamma_{2}}(V_{1}-V_{2})$. Thus, if $u:=v\gamma_{1}^{-1/2}$, then $\nabla\cdot(\gamma_{1}\nabla u)=\sqrt{\gamma_{1}\gamma_{2}}(V_{1}-V_{2}),$ with $u|_{\mathbb{S}^{d-1}}=1-\sqrt{\gamma_{2}/\gamma_{1}}$. By the uniform bounds in the ellipticity of $\gamma_{j}$, we have
$C^{-1}\left\lVert 1-\sqrt{\gamma_{2}/\gamma_{1}}\right\rVert_{H^{1}(\mathbb{B}^{d})}\leq\left\lVert V_{1}-V_{2}\right\rVert_{H^{-1}(\mathbb{B}^{d})}+\left\lVert 1-\sqrt{a_{2}/a_{1}}\right\rVert_{H^{1/2}(\partial\mathbb{B}^{d})}.$
Using that the trace at the boundary is constant, and defining $r_{d}:=\frac{2d}{d+2}$ for $d\geq 3$, $r_{2}:=(p+1)/2$, we obtain by the Sobolev embedding that
$C^{-1}\left\lVert 1-\sqrt{\gamma_{2}/\gamma_{1}}\right\rVert_{H^{1}(\mathbb{B}^{d})}\leq\left\lVert V_{1}-V_{2}\right\rVert_{L^{r_{d}}(\mathbb{B}^{d})}+|1-\sqrt{a_{2}/a_{1}}|,$
which yields
$C^{-1}\left\lVert\sqrt{\gamma}_{1}-\sqrt{\gamma_{2}}\right\rVert_{H^{1}(\mathbb{B}^{d})}\leq\left\lVert V_{1}-V_{2}\right\rVert_{L^{r_{d}}(\mathbb{B}^{d})}+|a_{1}-a_{2}|.$
Now, from (4.3) one gets
(A.1) $\left\lVert V_{j}\right\rVert_{L^{p}(\mathbb{B}^{d})}\leq C_{K}N.$
We use this and interpolation to go from $r_{d}$ to $1$ in the RHS:
(A.2) $C^{-1}\left\lVert\sqrt{\gamma}_{1}-\sqrt{\gamma_{2}}\right\rVert_{H^{1}(\mathbb{B}^{d})}\leq\left\lVert V_{1}-V_{2}\right\rVert_{L^{1}(\mathbb{B}^{d})}^{\beta_{d}}+|a_{1}-a_{2}|,$
where $\beta=\frac{p-r_{d}}{r_{d}(p-1)}$.
We now want to show that $\Delta v\in L^{r}(\mathbb{B}^{d})$ for some $1<r<p$. First we need to estimate the $vV_{1}$ term. If 1/q:=1/r1/p1/q:=1/r-1/p
$\left\lVert vV_{1}\right\rVert_{L^{r}}\leq\left\lVert V_{1}\right\rVert_{L^{p}(\mathbb{B}^{d})}\left\lVert v\right\rVert_{L^{q}(\mathbb{B}^{d})}\leq C\left\lVert V_{1}\right\rVert_{L^{p}(\mathbb{B}^{d})}\left\lVert v\right\rVert_{H^{1}(\mathbb{B}^{d})},$
by the Sobolev embedding, provided we can choose 1/q>1/21/d1/q>1/2-1/d. This gives us the condition 1/r1/p>1/21/d1/r-1/p>1/2-1/d, which is satisfied for any $1<r<p$ if $d=2$, or for $1<r\leq 2d/(2+d)$ when $d\geq 3$. In particular we can choose again $r=r_{d}$, which yields
$\left\lVert\Delta(\sqrt{\gamma}_{1}-\sqrt{\gamma_{2}})\right\rVert_{L^{r_{d}}(\mathbb{B}^{d})}\leq\left\lVert V_{1}\right\rVert_{L^{p}(\mathbb{B}^{d})}\left\lVert\sqrt{\gamma}_{1}-\sqrt{\gamma_{2}}\right\rVert_{H^{1}(\mathbb{B}^{d})}+\sqrt{K}\left\lVert V_{1}-V_{2}\right\rVert_{L^{r_{d}}(\mathbb{B}^{d})}.$
so the same interpolation as before in the RHS gives
$C^{-1}\left\lVert\Delta(\sqrt{\gamma}_{1}-\sqrt{\gamma_{2}})\right\rVert_{L^{r_{d}}(\mathbb{B}^{d})}\leq\left\lVert\sqrt{\gamma}_{1}-\sqrt{\gamma_{2}}\right\rVert_{H^{1}(\mathbb{B}^{d})}+\left\lVert V_{1}-V_{2}\right\rVert_{L^{1}(\mathbb{B}^{d})}^{\beta}.$
Since $r_{d}>1$, by Calderón-Zygmund estimates we have
$\left\lVert\sqrt{\gamma}_{1}-\sqrt{\gamma_{2}}-(\sqrt{a_{1}}-\sqrt{a_{2}})\right\rVert_{W^{2,r_{d}}(\mathbb{B}^{d})}\leq C\left\lVert\Delta(\sqrt{\gamma}_{1}-\sqrt{\gamma_{2}})\right\rVert_{L^{r_{d}}(\mathbb{B}^{d})},$
where we have subtracted the constant function $\sqrt{a_{1}}-\sqrt{a_{2}}$ to get a function with zero Dirichlet datum in the norm in the LHS. Putting the previous two estimates together, it follows that
$C^{-1}\left\lVert\sqrt{\gamma}_{1}-\sqrt{\gamma_{2}}\right\rVert_{W^{2,r_{d}}(\mathbb{B}^{d})}\leq\left\lVert\sqrt{\gamma}_{1}-\sqrt{\gamma_{2}}\right\rVert_{H^{1}(\mathbb{B}^{d})}+\left\lVert V_{1}-V_{2}\right\rVert_{L^{1}(\mathbb{B}^{d})}^{\beta}+|a_{1}-a_{2}|.$
Combining this estimate with (A.2) gives
$C^{-1}\left(\left\lVert\sqrt{\gamma}_{1}-\sqrt{\gamma_{2}}\right\rVert_{H^{1}(\mathbb{B}^{d})}+\left\lVert\sqrt{\gamma}_{1}-\sqrt{\gamma_{2}}\right\rVert_{W^{2,1}(\mathbb{B}^{d})}\right)\leq\left\lVert V_{1}-V_{2}\right\rVert_{L^{1}(\mathbb{B}^{d})}^{\beta}+|a_{1}-a_{2}|.$
To finish, we use that there exists a positive constant $C=C_{M,p,d}$ such that
(A.3) $\left\lVert{\gamma}_{1}-{\gamma_{2}}\right\rVert_{H^{1}(\mathbb{B}^{d})}+\left\lVert{\gamma}_{1}-{\gamma_{2}}\right\rVert_{W^{2,1}(\mathbb{B}^{d})}\\ \leq C\left(\left\lVert\sqrt{\gamma}_{1}-\sqrt{\gamma_{2}}\right\rVert_{H^{1}(\mathbb{B}^{d})}+\left\lVert\sqrt{\gamma}_{1}-\sqrt{\gamma_{2}}\right\rVert_{W^{2,1}(\mathbb{B}^{d})}\right).$
This follows from the following two facts. First, $\gamma_{1}-\gamma_{2}=\psi(\sqrt{\gamma}_{1}-\sqrt{\gamma_{2}})$, where $\psi:=\sqrt{\gamma}_{1}+\sqrt{\gamma_{2}}$ belongs to $W^{2,p}(\mathbb{B}^{d})$. Second, if $U$ is an open bounded set, one can verify that multiplication by a function in $W^{2,p}(U)$ with $p>d/2$ is bounded in both $W^{2,1}(U)$ and $H^{1}(U)$.
Local estimate: Let $I_{s}:=[s,1]$. If we define the radial profiles $u_{j}(r)$ and $q_{j}(r)$ by the identities $q_{j}(|x|)=V_{j}(x)$ and $u_{j}(|x|)=\sqrt{\gamma_{j}(x)}$, we have that
$r^{-(d-1)}\partial_{r}(r^{d-1}\partial_{r}u_{j}(r))+q_{j}(r)u_{j}(r)=0.$
If $w:=u_{1}-u_{2}$ it follows that
$(r^{d-1}w^{\prime})^{\prime}=r^{d-1}g,\qquad g:=-q_{1}w-(q_{1}-q_{2})u_{2}.$
Therefore
(A.4) $w^{\prime}(r)=r^{1-d}\left(w^{\prime}(1)-\int_{r}^{1}t^{d-1}g(t)\,dt\right),\qquad w(r)=w(1)-\int_{r}^{1}w^{\prime}(t)\,dt.$
Thus
$\left\lVert w^{\prime}\right\rVert_{L^{1}(I_{s})}\leq C_{d,s}\left(|w^{\prime}(1)|+\left\lVert g\right\rVert_{L^{1}(I_{s})}\right),\qquad\left\lVert w^{\prime\prime}\right\rVert_{L^{1}(I_{s})}\leq C_{d,s}\left(|w^{\prime}(1)|+\left\lVert g\right\rVert_{L^{1}(I_{s})}\right),$
since $w^{\prime\prime}(r)=\frac{d-1}{r}w^{\prime}-g$. a As a consequence, we have
$\left\lVert g\right\rVert_{L^{1}(I_{s})}\leq\left\lVert w\right\rVert_{L^{\infty}(0,1)}\left\lVert q_{1}\right\rVert_{L^{1}(I_{s})}+\sqrt{K}\left\lVert q_{1}-q_{2}\right\rVert_{L^{1}(I_{s})}.$
It remains to estimate $\left\lVert w\right\rVert_{L^{\infty}(I_{s})}$, which can be achieved using Grönwall’s inequality. Combining the identities in (A.4) and changing the order of integration, one obtains
$w(r)=w(1)+w^{\prime}(1)h(r)+\int_{r}^{1}J(r,t)(q_{1}(t)w(t)+(q_{1}(t)-q_{2}(t))u_{2}(t))\,dt.$
Since $h$ and $J$ are uniformly bounded, respectively, in $I_{s}$ and $I_{s}\times I_{s}$, one gets
$C^{-1}|w(r)|\leq|w(1)|+|w^{\prime}(1)|+\left\lVert q_{1}-q_{2}\right\rVert_{L^{1}(I_{s})}+\int_{r}^{1}|w(t)||q_{1}(t)|\,dt.$
By Grönwall’s inequality, it follows that
$C^{-1}|w(r)|\leq\left(|w(1)|+|w^{\prime}(1)|+\left\lVert q_{1}-q_{2}\right\rVert_{L^{1}(I_{s})}\right)\exp\left(\left\lVert q_{1}\right\rVert_{L^{1}(I_{s})}\right),$
so that
$C^{-1}\left\lVert w\right\rVert_{L^{\infty}}\leq|w(1)|+|w^{\prime}(1)|+\left\lVert q_{1}-q_{2}\right\rVert_{L^{1}(I_{s})}.$
Combining the previous estimates, we have shown that
$C^{-1}\left(\left\lVert w\right\rVert_{L^{1}(I_{s})}+\left\lVert w^{\prime}\right\rVert_{L^{1}(I_{s})}+\left\lVert w^{\prime\prime}\right\rVert_{L^{1}(I_{s})}\right)\leq|w(1)|+|w^{\prime}(1)|+\left\lVert q_{1}-q_{2}\right\rVert_{L^{1}(I_{s})}.$
To finish, one can use the same argument as in (A.3).